### v1

# reference.bib

Abstract
Optimization for robot control tasks, spanning various methodologies, includes Model Predictive Control (MPC). However, the complexity of the system, such as non-convex and non-differentiable cost functions and prolonged planning horizons often drastically increases the computation time, limiting MPC’s real-world applicability. Prior works in speeding up the optimization have limitations on solving convex problem and generalizing to hold out domains. To overcome this challenge, we develop a novel framework aiming at expediting optimization processes. In our framework, we combine offline self-supervised learning and online fine-tuning through reinforcement learning to improve the control performance and reduce optimization time.We demonstrate the effectiveness of our method on a novel, challenging Formula-1-track driving task, achieving $3.9\%$ higher performance in optimization time and $3.6\%$ higher performance in tracking accuracy on challenging holdout tracks.
I INTRODUCTION
Iterative control optimization algorithms have been widely adopted to control the dynamic systems such as autonomous vehicles [borrelli2005mpc, karnchanachari2020practical, kong2015kinematic, kim2022smooth], aircraft [bauersfeld2021mpc, jadbabaie2002control], humanoid robots [kuindersma2016optimization], etc. These optimization paradigms often entail dealing with complex systems featuring constraints, high dimensional solution space, and sudden state changes. Due to the inherent complexity of such systems, finding a good-enough solution (within some tolerance) in a single attempt can be exceedingly challenging for the optimizer. Consequently, solvers adopt an iterative approach to gradually converge towards a satisfactory solution. Typically, these algorithms start with an initial guess of the control inputs (e.g., joint accelerations) and employs solvers to iteratively minimize the cost function [sacks2023learning]. However, the optimization process can be significantly complicated by factors, such as the non-convex and non-differentiable cost function [rawlings1994nonlinear, eren2017model]. Consequently, the optimization time of MPC is typically a bottleneck in its real-world applications [bouzidi2023learning, lembono2020learning, mansard2018using, richter2009real]. Addressing this challenge is paramount for enhancing the practical utility of MPC in various domains.
Figure 1: Overview of our proposed algorithm. The first two blocks denote the two-phase training framework. In the first phase, we collect expert MPC demonstrations and train a warm-start policy using behavior cloning to speed up MPC. In the second phase, we fine-tune this policy within an online training framework to enhance its performance and generalizability. During testing, the proposed framework is evaluated on both training tracks and challenging holdout tracks, as demonstrated in the third block.
Given the iterative nature of the solvers used in optimization for robot control tasks, providing the solver with a better initial guess can expedite the optimization process. This process, known as warm starting, involves initializing the optimization algorithm with a solution that is closer to the optimal solution than a randomly chosen starting point. The intuition behind warm starting is that, at each iteration, the solver refines its solution based on previous iterations. As shown in Figure 2, a closer initial guess reduces the distance the optimization algorithm needs to traverse to converge to the optimal solution. By starting closer to the optimum, the solver can often avoid lengthy exploration of the solution space, leading to faster convergence and reduced computational effort. In this paper, we choose to test our proposed algorithm to warm start Model Predictive Control (MPC), which optimizes a trajectory over a predefined horizon.
To expedite the optimization process of MPC, various approaches have been adopted. However, they have fallen short on holdout domains, or have limitations in dealing with sudden state changes. Traditionally, one common technique involves utilizing the MPC solution from the previous sampling instance as the initial guess for the current control step [zeilinger2011real, pan2020imitation]. However, this approach falls short when faced with sudden state changes (e.g., the vehicle approaches a sharp turn). Another method involves maintaining a memory buffer to store historical MPC solutions, from which a suitable initial guess can be retrieved for future planning steps [mansard2018using, marcucci2020warm]. However, this approach may lack zero-shot generalizability. As the system’s state becomes heterogeneous, the size of the memory buffer needs to increase to maintain effectiveness. Moreover, searching for the initial guess within the memory buffer scales at least linearly with its size, potentially causing time-consuming operations. Alternatively, a learning-based methods have been proposed by Klaučo et al. to utilize k-NN classifier to classify the solution space into different active sets from which the solver searches for a solution [klauvco2019machine]. However, this method falls short in providing a precise warm-started initial guess and is not capable of dealing with challenging control tasks featuring long planning horizon and heterogeneous observations. Additionally, this approach is limited to solving strictly convex quadratic programs, imposing significant constraints on its applicability and generalizability given that many real-world control problems lack a linear or differentiable objective function and dynamics model. Thus, there is a pressing need for innovative approaches that can effectively address the challenges posed by the inherent complexity and non-linearity of real-world control systems.
In this paper, we propose a two-phase learning framework to learn a warm-start policy that provides a better initial guess for iterative control optimization algorithms (such as MPC) to reduce the optimization time. Our proposed algorithm is depicted in Figure 1, which includes a two-phase training framework and an additional holdout testing phase. An advantage of utilizing a learned policy to initialize the MPC, as opposed to learning a policy for end-to-end control of the system, is that it maintains the integrity of the original problem definition. The learned initial guess solely influences the starting point of the search, without altering the underlying problem setup, while MPC acts as a shield for the learned policy so that the control solution satisfies the constraints of the system, which ensures safer operation of the control system. Moreover, the proposed algorithm enables seamless adaptation to new conditions without the need for extensive retraining. During testing, the warm-start policy is tested on both the training domains and some holdout domains, demonstrating the generalizability of our framework in providing a good initial guess. Our key contributions are:
  1. 1.
    Propose a two-phase learning framework featuring offline training and online fine-tuning to learn a warm-start policy that provides the iterative control optimization algorithm with higher-quality initial guesses, aiding the command of a high-speed vehicle on multiple novel, challenging Formula 1 tracks in real-time where traditional warm start methods struggle.
  2. 2.
    Empirically evaluate our proposed two-phase learning framework and show the online fine-tuning phase helps the iterative control optimization algorithm achieve a $3.9\%$ higher performance in optimization time and a $3.6\%$ higher performance in tracking accuracy on challenging holdout tracks.
Figure 2: An illustrative demo showing how a better initial guess improves MPC optimization time.
II Preliminary
II-A Model Predictive Control
Unlike traditional control methods, MPC utilizes a predictive model of the system to anticipate future behavior of the system over a defined planning horizon, $H$. The control problem is formulated as an optimization task, where the objective is to minimize a cost function while satisfying constraints on the system’s inputs and outputs. This optimization problem is typically solved iteratively at each time step, generating a sequence of control inputs that steer the system towards a desired state while considering future predictions of its behavior. When MPC observes a new state, $x_{t}$, at time step, $t$, it optimizes a sequence of control inputs, $U$, over the planning horizon by iteratively minimizing a cost function $J$ while respecting system dynamics and constraints. Then, only the first action in the action sequence is applied to the controlled system, which leads the system to the next state. The optimization process is performed again at the next state. The MPC control law is formulated as shown below in Equations (1):
[…]
subject $\text{subject to}\quad x_{t+i+1}=f(x_{t+i},u_{t+i})$ (1b)
$U=[u_{t},...,u_{t+H-1}]\in\mathcal{U}_{j}\ \text{for all}\ j=1,\dots,n_{c_{u}}$ (1c)
$X=[x_{t},...,x_{t+H}]\in\mathcal{X}_{j}\ \text{for all}\ j=1,\dots,n_{c_{x}}$ (1d)
Equation (1a) formulates the objective function of the MPC. $H$ denotes the planning horizon, $t$ is the current step, and, $x_{t+i}$ is the system state at the $t+i$ step. Given the state , $x_{t+i}$ and the control input , $u_{t+i}$, the system dynamics model predicts the next state , $x_{t+i+1}$, using the dynamics model  (1b). Equation  (1c) and Equation  (1d) are the constraints for the control inputs and states. $n_{c_{u}}$ and $n_{c_{x}}$ are the number of constraints on control inputs and states respectively.
Solvers used to address the above optimization problem can be divided into gradient-based and gradient-free solvers. Gradient-based solvers leverage the gradient of the cost function with respect to the control inputs to iteratively adjust the control inputs towards the optimal solution [lee2011model, schwenzer2021review]. These solvers are effective when the cost function and constraints are smooth and differentiable, while gradient-free solvers do not rely on gradient information for optimization. Instead, gradient-free solvers explore the search space using heuristics, pattern searches, or stochastic techniques to find the optimal solution [lee2011model, schwenzer2021review]. To relax the constraints imposed on the problem formulation and demonstrate the generalizability of our framework, we choose to use the gradient-free solver COBYLA [powell1994direct] in our experiments.
III Method
In this section, we discuss in detail the proposed two-phase training framework. In the first phase, we run the MPC to collect the expert demonstrations, which are represented as state-action pairs. Then, we use behavior cloning to train a warm-start policy to mimic the expert MPC’s solution, as shown in Algorithm 1. The output of warm-start policy is utilized as an initial guess to warm start the MPC and reduce the optimization time. In the second phase, we load the pre-trained trajectory prediction model into an online training framework and fine tune the warm-start policy to address the suboptimality problem caused by behavior cloning and improve the model’s generalizability. The online fine tuning phase is shown in Algorithm 2.
Algorithm 1 Offline Training
1:Input: expert MPC $\pi^{\text{MPC}}$ with planning horizon $H$ and maximum optimization iteration $N_{expert}$ and all-zero vector $\vec{0}$ as initial guess, environment transition $T$, number of state-action pairs to collect $N$
2:Initialize neural network policy $\pi_{\theta}^{\text{warm}}$
3:$t\leftarrow 0,{\color[rgb]{0,0,0}s\leftarrow s_{0}},\mathcal{D}=\emptyset$
4:while  $t<N$ do
5:     $(u_{t},u_{t+1},\cdots,u_{t+H-1})\leftarrow\pi^{\text{MPC}}(s,\vec{0})$
6:     $\mathcal{D}\leftarrow\mathcal{D}\cup\{(s,u_{t},u_{t+1},\cdots,u_{t+H-1})\}$
7:     $s\leftarrow T(s,u_{t})$
8:     $t\leftarrow t+1$
9:end while
10:Train $\pi_{\theta}^{warm}$ with Equation (2) and $\mathcal{D}$
Algorithm 2 Online Fine Tuning
1:Input: fast MPC $\pi^{\text{MPC}}_{\text{fast}}$ with planning horizon $H$ and maximum optimization iteration $N_{fast}$, environment transition $T$, pre-trained warm-start policy $\pi_{\theta}^{warm}$
2:for each RL training iteration do
3:     Perceive an observation $s$
4:     $\hat{U}_{t}=(\hat{u}_{t},\hat{u}_{t+1},...,\hat{u}_{t+H-1})\leftarrow\pi^{warm}(s)$
5:     $pos_{i+1}^{car}=M_{dynamics}((s_{i},\hat{u}_{i}))|_{i=t}^{H-1}$
6:     Compute accumulated $xte$ using Equation (4)
7:     $U_{t}^{\text{MPC}}=(u_{t},u_{t+1},...,u_{t+H-1})\leftarrow\pi^{\text{MPC}}_{\text{fast}}(s,\hat{U}_{t})$
8:     Calculate reward in RL using Equation (5)
9:     $L_{imitation}=MSE(U_{t}^{MPC},U_{t}^{guess})$
10:     Compute training loss $L$ in Equation (3)
11:     $s\leftarrow T(s,u_{t})$
12:     Update $\pi_{\theta}^{warm}$ with $L$
13:end for
III-A Offline Training
At this phase, we implement an expert MPC to collect a dataset $\mathcal{D}$ containing $N$ state-action pairs first. The expert MPC, $\pi^{MPC}$ aims to control the agent to finish the task without relying on a warm-started initial guess for the control action sequence. At each step, $\pi^{MPC}$ observes a state $s$ and takes an all-zero vector as the initial guess and optimizes that initial guess to output $(u_{t},u_{t+1},\cdots,u_{t+H-1})$ (line 5). Then the state-action pair, $(s,u_{t},u_{t+1},\cdots,u_{t+H-1})$ is stored in the dataset $\mathcal{D}$ (line 6). The first action in the action sequence is applied to the system, which leads the system to the next state based on the environment transition $T$ (line 7-8). During the data collection stage, $\pi^{MPC}$ optimizes the initial guess for enough iterations to make sure it achieves a high control performance by disabling the “early stop” for the expert MPC.
We design our warm-start policy, $\pi_{\theta}^{warm}$, as a multi-layer perceptron with ReLU activation function [glorot2010understanding]. Given the current state of the vehicle, the $\pi_{\theta}^{warm}$ predicts a sequence of actions which then serves as the initial guess of the MPC to warm start the optimization process. In the offline training phase, we utilize behavior cloning to train $\pi_{\theta}^{warm}$ (line 10).
Behavior Cloning (BC) is a simple yet effective algorithm to learn from demonstrations. The demonstration is a set of trajectories: $\mathcal{D}=\{\tau_{i}\}$. BC learns a control policy, $\pi_{\theta}$, by minimizing the Mean Squared Error (MSE) in the demonstration set, shown in Equation (2). Our algorithm utilizes pre-collected MPC data to perform BC as a warm start for the MPC to improves the MPC’s optimization time.
$\theta^{*}=\min_{\theta}{\sum_{\tau^{i}\in\mathcal{D}}\sum_{t=0}^{T}{(\pi_{\theta}^{warm}(s^{i})_{t}-u_{t}^{i})^{2}}}$ (2)
III-B Online Fine Tuning
Our second phase, online fine tuning, further improves the warm-start policy by learning from the data gathered online using the policy trained through behavior cloning. This phase combines the best aspects of Reinforcement Learning and Dataset Aggregation (DAgger) [ross2011reduction] algorithm. Reinforcement Learning (RL) operates under the formalization of Markov Decision Process (MDP), $\mathcal{M}=\langle\mathcal{S},\mathcal{A},R,T,\gamma,\rho_{0}\rangle$. $\mathcal{S}$ is the state space and $\mathcal{A}$ denotes the action space. $R$ encodes the reward of a given state. $T$ is a deterministic transition function that decides the next state, $s^{\prime}$, when applying the action, $a\in A$, in state, $s\in S$. $\gamma\in(0,1)$ is the temporal discount factor. $\rho_{0}$ denotes the initial state probability distribution. A policy, $\pi:S\rightarrow[0,1]^{|A|}$, is a mapping from states to actions or to a probability distribution over actions. The objective of RL is to find the policy that optimizes the expected discounted return, $\pi^{*}=\mathbb{E}_{\tau\sim\pi}\left[\sum^{\infty}_{t=0}\gamma^{t}R(s_{t})\right]$.
DAgger builds upon behavior cloning (BC) by incorporating online interaction with the environment and online querying of the expert. Unlike BC, which trains solely on a fixed dataset of expert demonstrations, DAgger actively collects data from interactions with the environment and solicits expert feedback to augment its training. This online improvement process allows DAgger to learn from a more diverse set of experiences, adapt to new situations, and refine the agent’s policy over time, ultimately leading to improved performance in imitation learning tasks.
In our framework, RL is used to address the sub-optimality and covariance shift problem caused by offline behavior cloning. DAgger, on the one hand, is adopted to regulate the trajectory prediction model during reinforcement learning training, ensuring it does not forget the experience learned from behavior cloning. Further more, DAgger improves the performance of imitation learning by expanding the diversity of the expert MPC’s demonstrations. Given that offline reinforcement learning algorithms require lots of data to effectively cover the entire state space, we choose to fine-tune our pre-trained trajectory prediction model using Proximal Policy Optimization (PPO) [schulman2017proximal] as our RL training algorithm. The trajectory prediction model acts as a policy network in the PPO framework. DAgger is integrated as a loss term in the RL training, as shown in Equation (3). $L_{RL}$ is the standard PPO loss which includes three parts: $L_{policy}$, $L_{entropy}$, and $L_{value}$. $L_{imitation}$ is the loss signal from DAgger. It represents MSE between the expert MPC’s solution and the warm-started initial guess output from $\pi_{\theta}^{warm}$. $\lambda$ is the weight coefficient for the RL loss and imitation loss.
$L$ $=\lambda\cdot L_{RL}+(1-\lambda)\cdot L_{imitation}$ (3)
$L_{RL}$ $=\lambda_{1}\cdot L_{policy}+\lambda_{2}\cdot L_{entropy}+\lambda_{3}\cdot L_{value}$
$xte=distance(pos^{car}_{t},\mathbf{p}_{i}(x_{i},y_{i})^{closest})$ (4)
In our driving task, the reward at each step of the RL training is shown in Equation (5). The first term is the negative MPC optimization time, and the second term is the negative of the accumulated Cross Track Error ($xte$) over the planning horizon, $H$. The $xte$ is defined in Equation (4). At each step, $\pi_{\theta}^{warm}$ will output an action sequence of length $H$ based on the current state vector, $s_{t}$. Then, the dynamics model, $M_{dynamics}$, is used to calculate the future positions of the agent. The $xte$ is computed between the reference trajectory and the future positions of the vehicle, $pos_{i}^{car}$. The reward obtained at each step directly influences $L_{RL}$ by contributing to the calculation of the advantage estimate, which measures the discrepancy between the observed reward and the expected value of the state-action pair. This estimate affects both $L_{policy}$, which encourages actions leading to higher rewards, and $L_{value}$, which trains the value function to better estimate cumulative rewards.
$r_{t}=-time_{MPC}-\sum_{i=t+1}^{t+H}xte(pos_{i}^{car},Traj^{ref})$ (5)
This reward design helps to optimize the $\pi_{\theta}^{warm}$ by minimizing the MPC running time and by minimizing the difference between the planned trajectory and the reference trajectory, thereby reducing MPC optimization time. This helps to address the sub-optimality problem caused by behavior cloning.
IV EXPERIMENTS
In this section, we will discuss in detail the setup of the Formula 1 tracks domains, followed by the MPC design in this testing domain. Additionally, we will present experimental results from both training and holdout tracks, demonstrating how our algorithm enhances MPC’s performance in terms of both optimization time and tracking accuracy.
IV-A Experiment Setup
The experiments are done in high-speed Formula 1 tracks [f1tenth_racetracks] which are divided into three training tracks and seven zero-shot testing tracks, as shown in Figure 3(a), Figure 3(b), and Figure 3(c), respectively. The reference trajectory is represented as a set of waypoints in the center of the track. The track size is downscaled to a 10:1 ratio to make each lap a reasonable length. The reference speed for the vehicle in the expert MPC is $10m/s$. Given the downscaled size of the track data, the vehicle could actually speed up to $100m/s$ in the 1:1 scale track. The friction between the tire and the road is not considered in our dynamics model, as it depends on multiple factors such as tire pressure, temperature, humidity, and road conditions. The path tracking is challenging for traditional gradient-free MPC because the vehicle is running at a high speed and the tracks have multiple sharp turns. In this scenario, traditional gradient-free MPC solvers could not optimize the control solution in real time, which makes it unreasonable for real-world application.
The three tracks used for demonstration collection and training are shown in Figure 3(a). The six zero-shot tracks are not presented to the vehicle before testing to demonstrate our algorithm’s zero-shot generalizability. During training, the expert MPC is first rolled out on training tracks to collect demonstrations for offline training introduced in Section III-A . Then, the warm-start policy is fine-tuned on the same three tracks using the algorithm introduced in Section III-B.
As shown in Figure 2, without a specified threshold, optimization continues until reaching the maximum iteration limit. In our experiments, we found that the MPC cost reduces subtly during the later optimization iterations. As such, it is important to provide the solver with an early-stop criterion. We implement the early-stop condition based on a planned trajectory’s accumulated cross-track error ($xte$) being $<0.1m$. Additionally, to ensure real-time planning during testing, we cap the maximum optimization iterations at 50, which guarantees that the optimization time is less than 0.08 seconds in worst-case scenarios. Details of the MPC design and hyper parameters in MPC and training are presented in Appendix VI-A and VI-B
(a) Training tracks
(b) Zero-shot tracks (complex)
(c) Zero-shot tracks (simple)
Figure 3: Tracks tested during training and testing.
IV-B Experiment Results
During testing, we evaluate on two metrics: the average MPC optimization time (seconds) per step and the average $xte$ (m) per step. We benchmark our algorithm’s performance against MPC using all-zero initial guesses and initial guesses derived from the MPC’s solution at the previous step. We perform testing on both the training tracks (Figure 3(a)) and the challenging zero-shot tracks (Figure 3(b)). The results are shown in Figure 4 and Figure 5, respectively.
IV-B1 Optimization Time
On both the training and zero-shot tracks, employing a warm-start policy trained through either offline learning or a combination of offline and online fine-tuning significantly reduces MPC optimization time, falling well within the upper bound of $0.08s$. Additionally, the warm-start policy trained with both offline and online fine-tuning achieves better MPC optimization time compared to the policy trained solely through offline behavior cloning, demonstrating the capability of our online fine-tuning algorithm in addressing the suboptimality problem caused by behavior cloning (BC). Furthermore, our two-phase training framework achieves more significant improvement in the zero-shot domains, indicating that the online fine-tuning phase alleviates the covariance shift problem caused by BC and enhances the model’s generalizability.
IV-B2 Control Performance ($xte$)
In the context of the training tracks, the warm-start policy trained through offline and online fine-tuning outperforms the policy trained solely through offline behavior cloning in terms of tracking accuracy, further illustrating the effectiveness of our online fine-tuning algorithm in addressing the suboptimality problem caused by BC. Given that the width of the downscaled track is $2.2m$, achieving an $xte$ of less than $0.3m$ demonstrates precise tracking of the reference path. Additionally, our two-phase training framework exhibits more significant improvement in the zero-shot domains, reinforcing the efficacy of the online fine-tuning phase in mitigating the covariance shift problem caused by BC and improving the model’s generalizability.
Figure 4: Experiment results on training tracks.
Figure 5: Experiment results on the six zero-shot testing tracks.
The results also reveal that employing an MPC with either an all-zero initial guess or an initial guess derived from the MPC’s previous solution fails to complete laps on both the training and zero-shot tracks. The vehicle consistently deviates from the lane and struggles to navigate sharp turns. This limitation arises because the MPC often requires more iterations to optimize the control solution as the vehicle approaches the curves of the track. Since the real-time MPC only optimizes the solution for a maximum of 50 iterations at each step, the returned solution lacks the optimization necessary to guide the vehicle through sharp turns effectively. This underscores the necessity of a well-informed initial guess to minimize the number of optimization iterations required.
As the real-time MPC with either an all-zero initial guess or an initial guess derived from the MPC’s previous solution fails to complete laps on both the training and challenging zero-shot tracks, we conducted further testing on four warm start policies on a simple zero-shot track, IMS, as depicted in Figure 3(c). The results are summarized in Figure 6. Our method demonstrates a significant reduction in MPC optimization time while maintaining high tracking accuracy, showcasing its effectiveness in enhancing MPC performance.
Figure 6: Experiment results on the simple IMS track.
As the vehicle route on the track, we further compute the curvature of the track surrounding the vehicle. This analysis allows us to illustrate the correlation between the curvature and lateral error ($xte$), depicted in Figure 7. It shows that our second phase fine tuning has an advantage on average $xte$ and this advantage has a modest increase when the track is less curvy. This outcome underscores the efficacy of our proposed algorithm in enhancing the precision of the planned trajectory. Details regarding the method employed to compute the curvature of the track around the vehicle are provided in the Appendix VI-C
Figure 7: We run the MPC with two types of warm-start policy (BC and ours) on the ”Catalunya” track, where we gather data on the $xte$ and track curvature surrounding the vehicle. Subsequently, we depict these data points in the figure above.
In summary, our empirical results support that:
  • MPC with initial guesses trained by data-driven methods significantly outperforms those with all-zero or previous solution-derived guesses, crucial for navigating sharp turns.
  • The warm-start policy trained with our two-phase learning framework reduces MPC optimization time and $xte$ on both training and zero-shot tracks, showing the capability of our online fine-tuning algorithm in addressing the suboptimality problem.
  • The two-phase training shows superior performance in zero-shot scenarios, indicating better generalizability.
V LIMITATIONS AND FUTURE WORKS
While our proposed two-phase learning framework shows promising results in expediting optimization processes and enhancing control performance for robot control tasks, there are several limitations and avenues for future research that merit consideration. One limitation of our current experiments is that we have only focused on autonomous driving where the action space is limited to two dimensions (i.e., steering and acceleration). Future work could expand into more complex domains, such as robot arm manipulation or quadcopter control, to explore the algorithm’s efficacy in higher-dimensional spaces. Additionally, future research could also involve utilizing our proposed algorithm to expedite other iterative control optimization techniques like MPPI. Although our algorithm primarily interfaces with MPC, it is versatile enough to serve as a warm start for other iterative control optimization methods like Model Predictive Path Integral (MPPI) control [williams2015model]. MPPI typically employs stochastic optimization, initializing with a guess of the control sequence distribution, sampling sequences, and iteratively refining the distribution to minimize control cost. Our framework can be trained to predict the trajectory Gaussian distribution mean, enhancing MPPI’s performance by offering a superior initial distribution. This reduces the iteration count needed for optimization, aligning with early stopping criteria. While MPC suits systems with complex dynamics and constraints, we choose it as our focus due to its representativeness in trajectory optimization methods.
VI CONCLUSIONS
In this paper, we introduce a novel approach to accelerate MPC optimization by learning a warm-start policy. Our two-phase learning framework combines offline behavior cloning and online fine-tuning to provide improved initial guesses for the MPC solver. Experimental results on both training and zero-shot tracks demonstrate the effectiveness of our approach in reducing optimization time and enhancing path tracking accuracy. Our learning framework integrated with MPC opens up new avenues for improving the efficiency and applicability of trajectory optimization in various dynamic systems.
APPENDIX
VI-A MPC Design Details
MPC is designed to control the acceleration and the steering of the vehicle to track the reference trajectory. The dynamics model of the vehicle, $M_{dynamics}$, is shown in Equation (6). $(x_{t}^{car},y_{t}^{car})$ is the global position of the vehicle at time step $t$. $v_{t}$ and $yaw_{t}$ are the speed ($m/s$) and yaw angle ($rad$) of the vehicle at time step $t$. $a_{t}$ and $\theta_{t}^{\text{steering}}$ are acceleration ($m/s^{2}$) and steering angle ($rad$) inputs to the vehicle at time step $t$. $L$ is the wheelbase of the vehicle. Here we use Ford Mustang’s dimensions, $L=2.89m$. And $dt=0.02$ is the length of each time step which represents the duration between successive updates of control inputs and system states.
$x_{t+1}^{car}=x_{t}^{car}+v_{t}\cdot\cos(yaw_{t})\cdot dt$ (6a)
$y_{t+1}^{car}=y_{t}^{car}+v_{t}\cdot\sin(yaw_{t})\cdot dt$ (6b)
$yaw_{t+1}=yaw_{t}+\frac{v_{t}}{L}\cdot\tan(\theta_{t}^{\text{steering}})\cdot dt$ (6c)
$v_{t+1}=v_{t}+a_{t}\cdot dt$ (6d)
The MPC objective function is composed of five parts as shown in Equation (7). The first two terms are $xte$ and Error in Heading ($eth$). $xte$ is computed using Equation 4. $eth$ denotes the angular disparity between the intended path direction and the current heading of a vehicle in path tracking systems. And $eth$ is computed using Equation (8). $v_{t}^{\text{ref}}$ is the desired speed of the vehicle. The last two terms in Equation (7) regulate the rate of change of the steering angle and acceleration to make planned trajectory smoother. $w_{0},w_{1},w_{2},w_{3},w_{4}$ are the coefficients balancing the importance of each term. The planning horizon of the MPC is 25 steps and the planning step $dt$ is 0.02 seconds, which means that the MPC looks 0.5 seconds ahead.
$J_{t}$ $=\sum_{i=t}^{t+H-1}(w_{0}\cdot xte_{i}^{2}+w_{1}\cdot eth_{i}^{2}$ (7)
$\quad+w_{2}\cdot(v_{i}-v_{t}^{\text{ref}})^{2}$
$\quad+w_{3}\cdot(steer_{i}-steer_{i-1})^{2}$
$\quad+w_{4}\cdot(throttle_{i}-throttle_{i-1})^{2})$
$eth=abs\left(yaw^{car}_{t}-\arctan\left({\frac{y_{i+1}-y_{i}}{x_{i+1}-x_{i}}}\right)\right)$ (8)
VI-B Hyper Parameters in MPC and Training
Hyperparameters Values
Weight in MPC cost function $w_{0}$ 2000
Weight in MPC cost function $w_{1}$ 100
Weight in MPC cost function $w_{2}$ 60
Weight in MPC cost function $w_{3}$ 2
Weight in MPC cost function $w_{4}$ 20
Weight in MPC cost function $w_{5}$ 300
Loss weight coefficient $\lambda$ 0.9
Loss weight coefficient $\lambda_{1}$ 0.5
Loss weight coefficient $\lambda_{2}$ 0
Loss weight coefficient $\lambda_{3}$ 0.5
MPC planning horizon 25
Length of each time step 0.02 second
Reference speed $10m/s$
Wheelbase 2.89 m
Expert MPC maximum optimization iteration 300
Real-time MPC maximum optimization iteration 300
Real time MPC early stop criteria $xte<0.1$
VI-C Curvature Computation
Algorithm 3 Curvature Computation
1:Extract 10 consecutive nearest waypoints ahead of the vehicle, $\mathbf{P}=\{\mathbf{p}_{1},\mathbf{p}_{2},...,\mathbf{p}_{10}\}$. $\mathbf{p}_{i}=(x_{i},y_{i})$ represent the $i$-th waypoint
2:for $\mathbf{p}_{i-1}$, $\mathbf{p}_{i}$, and $\mathbf{p}_{i+1}$ in $\mathbf{P}$ do
3:     Compute the vector $\mathbf{v}_{1}=\mathbf{p}_{i}-\mathbf{p}_{i-1}$
4:     Compute the vector $\mathbf{v}_{2}=\mathbf{p}_{i+1}-\mathbf{p}_{i}$
5:     Calculate the dot product $\mathbf{v}_{1}\cdot\mathbf{v}_{2}$
6:     Calculate the magnitudes $\|\mathbf{v}_{1}\|$ and $\|\mathbf{v}_{2}\|$
7:     Compute the $\cos$ of the angle between the vectors:
$\cos(\theta_{i})=\left(\frac{\mathbf{v}_{1}\cdot\mathbf{v}_{2}}{\|\mathbf{v}_{1}\|\|\mathbf{v}_{2}\|}\right)$
8:end for
9:$Curvature=\sum_{i=2}^{9}\left(1-cos(\theta_{i})\right)$
[…]

### v2

@@
-reference.bib
+Faster Model Predictive Control via Self-Supervised Initialization Learning
 Abstract
@@
 However, the complexity of the system, such as non-convex and non-differentiable cost functions and prolonged planning horizons often drastically increases the computation time, limiting MPC’s real-world applicability.
-Prior works in speeding up the optimization have limitations on solving convex problem and generalizing to hold out domains.
-To overcome this challenge, we develop a novel framework aiming at expediting optimization processes.
-In our framework, we combine offline self-supervised learning and online fine-tuning through reinforcement learning to improve the control performance and reduce optimization time.We demonstrate the effectiveness of our method on a novel, challenging Formula-1-track driving task, achieving $3.9\%$ higher performance in optimization time and $3.6\%$ higher performance in tracking accuracy on challenging holdout tracks.
+Prior works in speeding up the optimization have limitations on optimizing MPC running time directly and generalizing to hold out domains.
+To overcome this challenge, we develop a novel framework aiming at expediting optimization processes directly.
+In our framework, we combine offline self-supervised learning and online fine-tuning to improve the control performance and reduce optimization time.
+We demonstrate the success of our method on a novel and challenging Formula 1 track driving task.
+Comparing to single-phase training, our approach achieves a $19.4\%$ reduction in optimization time and a $6.3\%$ improvement in tracking accuracy on zero-shot tracks.
 I INTRODUCTION
-Iterative control optimization algorithms have been widely adopted to control the dynamic systems such as autonomous vehicles [borrelli2005mpc, karnchanachari2020practical, kong2015kinematic, kim2022smooth], aircraft [bauersfeld2021mpc, jadbabaie2002control], humanoid robots [kuindersma2016optimization], etc.
-These optimization paradigms often entail dealing with complex systems featuring constraints, high dimensional solution space, and sudden state changes.
-Due to the inherent complexity of such systems, finding a good-enough solution (within some tolerance) in a single attempt can be exceedingly challenging for the optimizer.
-Consequently, solvers adopt an iterative approach to gradually converge towards a satisfactory solution.
-Typically, these algorithms start with an initial guess of the control inputs (e.g., joint accelerations) and employs solvers to iteratively minimize the cost function [sacks2023learning].
-However, the optimization process can be significantly complicated by factors, such as the non-convex and non-differentiable cost function [rawlings1994nonlinear, eren2017model].
-Consequently, the optimization time of MPC is typically a bottleneck in its real-world applications [bouzidi2023learning, lembono2020learning, mansard2018using, richter2009real].
+Algorithms that optimize control outputs iteratively like MPC have been widely adopted to control dynamic systems, such as autonomous vehicles [1, 2, 3, 4], aircraft [5, 6], humanoid robots [7], etc.
+However, MPC in real-world applications is challenging due to computational time bottlenecks [8, 9, 10, 11], particularly with complex systems featuring non-convex and non-differentiable cost functions and prolonged planning horizons.
+[12, 13].
 Addressing this challenge is paramount for enhancing the practical utility of MPC in various domains.
@@
 In the second phase, we fine-tune this policy within an online training framework to enhance its performance and generalizability.
-During testing, the proposed framework is evaluated on both training tracks and challenging holdout tracks, as demonstrated in the third block.
-Given the iterative nature of the solvers used in optimization for robot control tasks, providing the solver with a better initial guess can expedite the optimization process.
-This process, known as warm starting, involves initializing the optimization algorithm with a solution that is closer to the optimal solution than a randomly chosen starting point.
+During testing, the proposed framework is evaluated on both training tracks and challenging zero-shot tracks, as demonstrated in the third block.
+Warm starting in robot control optimization involves initializing the solver with a closer-to-optimal solution, speeding up the iterative optimization process.
 The intuition behind warm starting is that, at each iteration, the solver refines its solution based on previous iterations.
@@
 In this paper, we choose to test our proposed algorithm to warm start Model Predictive Control (MPC), which optimizes a trajectory over a predefined horizon.
-To expedite the optimization process of MPC, various approaches have been adopted.
-However, they have fallen short on holdout domains, or have limitations in dealing with sudden state changes.
-Traditionally, one common technique involves utilizing the MPC solution from the previous sampling instance as the initial guess for the current control step [zeilinger2011real, pan2020imitation].
+To provide MPC with a high quality initial guess, various approaches have been adopted.
+Traditionally, one common technique involves utilizing the MPC solution from the previous sampling instance as the initial guess for the current control step [14, 15].
 However, this approach falls short when faced with sudden state changes (e.g., the vehicle approaches a sharp turn).
-Another method involves maintaining a memory buffer to store historical MPC solutions, from which a suitable initial guess can be retrieved for future planning steps [mansard2018using, marcucci2020warm].
+Another method involves maintaining a memory buffer to store historical MPC solutions, from which a suitable initial guess can be retrieved for future planning steps [10, 16].
 However, this approach may lack zero-shot generalizability.
-As the system’s state becomes heterogeneous, the size of the memory buffer needs to increase to maintain effectiveness.
-Moreover, searching for the initial guess within the memory buffer scales at least linearly with its size, potentially causing time-consuming operations.
-Alternatively, a learning-based methods have been proposed by Klaučo et al.
-to utilize k-NN classifier to classify the solution space into different active sets from which the solver searches for a solution [klauvco2019machine].
-However, this method falls short in providing a precise warm-started initial guess and is not capable of dealing with challenging control tasks featuring long planning horizon and heterogeneous observations.
-Additionally, this approach is limited to solving strictly convex quadratic programs, imposing significant constraints on its applicability and generalizability given that many real-world control problems lack a linear or differentiable objective function and dynamics model.
+As the system’s state becomes heterogeneous, the size of the memory buffer needs to increase, thus causing time-consuming operations to searching for the initial guess within the buffer.
+Alternatively Klaučo et al.
+propose using a k-NN classifier to classify the solution space into active sets for the solver to search for a solution [17].
+However, this method lacks precise warm-start guesses and struggles with complex control tasks involving long planning horizons and heterogeneous observations.
+Moreover, it’s limited to strictly convex quadratic programs, reducing its applicability to real-world problems with non-linear or non-differentiable objectives and dynamics.
+More recent approaches have adopted supervised learning to boost the performance of model predictive optimization algorithms.
+[18, 19] adopts behavior cloning (BC) to learn the distribution of the initial guesses for the solver.
+However, supervised learning cannot directly learn to minimize the computation time of MPC.
+Instead, all supervised learning can do is to imitate the output of the solver.
+Besides, BC also suffers from covariance shift, which degrades model’s performance on zero-shot scenarios [20].
+[21] used reinforcement learning (RL) to train a warm-start policy to help quadrotor track simple zig-zag trajectories.
+Yet, RL algorithms are too data hungry to learn an optimal policy from scratch, posing challenge to complex, long-horizon tasks [22].
+Furthermore, [21] is still not directly learning to minimize the computation time of MPC.
 Thus, there is a pressing need for innovative approaches that can effectively address the challenges posed by the inherent complexity and non-linearity of real-world control systems.
 In this paper, we propose a two-phase learning framework to learn a warm-start policy that provides a better initial guess for iterative control optimization algorithms (such as MPC) to reduce the optimization time.
-Our proposed algorithm is depicted in Figure 1, which includes a two-phase training framework and an additional holdout testing phase.
-An advantage of utilizing a learned policy to initialize the MPC, as opposed to learning a policy for end-to-end control of the system, is that it maintains the integrity of the original problem definition.
-The learned initial guess solely influences the starting point of the search, without altering the underlying problem setup, while MPC acts as a shield for the learned policy so that the control solution satisfies the constraints of the system, which ensures safer operation of the control system.
-Moreover, the proposed algorithm enables seamless adaptation to new conditions without the need for extensive retraining.
-During testing, the warm-start policy is tested on both the training domains and some holdout domains, demonstrating the generalizability of our framework in providing a good initial guess.
+Our proposed algorithm is depicted in Figure 1, which includes a two-phase training framework and an additional zero-shot testing phase.
+The two training phase works reciprocally to each other.
+On the one hand, the online-fine tuning phase uses RL to address the suboptimality and covariance shift problem of the offline BC by directly optimize the computation time of MPC.
+On the other hand, the offline BC provides online learning algorithm with a good starting point to expedite the training process.
+Moreover, another advantage of using a learned policy to initialize the MPC, rather than for end-to-end control, is that it preserves the original definition of the control problem.
+The learned policy only affects the starting point, while MPC ensures the control solution satisfies system constraints, leading to safer operation.
 Our key contributions are:
   1. 1.
-    Propose a two-phase learning framework featuring offline training and online fine-tuning to learn a warm-start policy that provides the iterative control optimization algorithm with higher-quality initial guesses, aiding the command of a high-speed vehicle on multiple novel, challenging Formula 1 tracks in real-time where traditional warm start methods struggle.
+    Propose a two-phase learning framework featuring offline training and online fine-tuning to learn a warm-start policy that provides the iterative control optimization algorithm with higher-quality initial guesses, aiding the command of a high-speed vehicle on multiple novel, challenging Formula 1 tracks in real-time where traditional warm start methods all failed.
   2. 2.
-    Empirically evaluate our proposed two-phase learning framework and show the online fine-tuning phase helps the iterative control optimization algorithm achieve a $3.9\%$ higher performance in optimization time and a $3.6\%$ higher performance in tracking accuracy on challenging holdout tracks.
+    Empirically evaluate our proposed two-phase learning framework and show the online fine-tuning phase helps the iterative control optimization algorithm achieve a $19.4\%$ higher performance in optimization time and a $6.3\%$ higher performance in tracking accuracy on challenging zero-shot tracks comparing to single-phase training.
 Figure 2: An illustrative demo showing how a better initial guess improves MPC optimization time.
-II Preliminary
+II Preliminaries
 II-A Model Predictive Control
-Unlike traditional control methods, MPC utilizes a predictive model of the system to anticipate future behavior of the system over a defined planning horizon, $H$.
-The control problem is formulated as an optimization task, where the objective is to minimize a cost function while satisfying constraints on the system’s inputs and outputs.
+MPC is formulated as an optimization task, where the objective is to minimize a cost function while satisfying constraints on the system’s inputs and outputs over a finite horizon, H.
 This optimization problem is typically solved iteratively at each time step, generating a sequence of control inputs that steer the system towards a desired state while considering future predictions of its behavior.
-When MPC observes a new state, $x_{t}$, at time step, $t$, it optimizes a sequence of control inputs, $U$, over the planning horizon by iteratively minimizing a cost function $J$ while respecting system dynamics and constraints.
+When MPC observes a new state, $x_{t}$, at time step, $t$, it optimizes a sequence of control inputs, $U$, over the planning horizon by iteratively minimizing a cost function , $J$ while respecting system dynamics and constraints.
 Then, only the first action in the action sequence is applied to the controlled system, which leads the system to the next state.
@@
 $n_{c_{u}}$ and $n_{c_{x}}$ are the number of constraints on control inputs and states respectively.
-Solvers used to address the above optimization problem can be divided into gradient-based and gradient-free solvers.
-Gradient-based solvers leverage the gradient of the cost function with respect to the control inputs to iteratively adjust the control inputs towards the optimal solution [lee2011model, schwenzer2021review].
-These solvers are effective when the cost function and constraints are smooth and differentiable, while gradient-free solvers do not rely on gradient information for optimization.
-Instead, gradient-free solvers explore the search space using heuristics, pattern searches, or stochastic techniques to find the optimal solution [lee2011model, schwenzer2021review].
-To relax the constraints imposed on the problem formulation and demonstrate the generalizability of our framework, we choose to use the gradient-free solver COBYLA [powell1994direct] in our experiments.
 III Method
@@
 In the first phase, we run the MPC to collect the expert demonstrations, which are represented as state-action pairs.
-Then, we use behavior cloning to train a warm-start policy to mimic the expert MPC’s solution, as shown in Algorithm 1.
+Then, we use BC to train a warm-start policy to mimic the expert MPC’s solution, as shown in Algorithm 1.
 The output of warm-start policy is utilized as an initial guess to warm start the MPC and reduce the optimization time.
-In the second phase, we load the pre-trained trajectory prediction model into an online training framework and fine tune the warm-start policy to address the suboptimality problem caused by behavior cloning and improve the model’s generalizability.
+In the second phase, we load the pre-trained trajectory prediction model into an online training framework and fine tune the warm-start policy to address the suboptimality problem caused by BC and improve the model’s generalizability.
 The online fine tuning phase is shown in Algorithm 2.
@@
 2:Initialize neural network policy $\pi_{\theta}^{\text{warm}}$
-3:$t\leftarrow 0,{\color[rgb]{0,0,0}s\leftarrow s_{0}},\mathcal{D}=\emptyset$
+3:$t\leftarrow 0,s\leftarrow s_{0},\mathcal{D}=\emptyset$
 4:while $t<N$ do
@@
 The first action in the action sequence is applied to the system, which leads the system to the next state based on the environment transition $T$ (line 7-8).
-During the data collection stage, $\pi^{MPC}$ optimizes the initial guess for enough iterations to make sure it achieves a high control performance by disabling the “early stop” for the expert MPC.
-We design our warm-start policy, $\pi_{\theta}^{warm}$, as a multi-layer perceptron with ReLU activation function [glorot2010understanding].
+During the data collection stage, $\pi^{MPC}$ optimizes the initial guess for enough iterations to ensure it achieves a high control performance by disabling the “early stop” for the expert MPC.
+We design our warm-start policy, $\pi_{\theta}^{warm}$, as a multi-layer perceptron with ReLU activation function [23].
 Given the current state of the vehicle, the $\pi_{\theta}^{warm}$ predicts a sequence of actions which then serves as the initial guess of the MPC to warm start the optimization process.
-In the offline training phase, we utilize behavior cloning to train $\pi_{\theta}^{warm}$ (line 10).
-Behavior Cloning (BC) is a simple yet effective algorithm to learn from demonstrations.
+In the offline training phase, we utilize BC to train $\pi_{\theta}^{warm}$ (line 10).
+BC is a simple yet effective algorithm to learn from demonstrations.
 The demonstration is a set of trajectories: $\mathcal{D}=\{\tau_{i}\}$.
@@
 III-B Online Fine Tuning
-Our second phase, online fine tuning, further improves the warm-start policy by learning from the data gathered online using the policy trained through behavior cloning.
-This phase combines the best aspects of Reinforcement Learning and Dataset Aggregation (DAgger) [ross2011reduction] algorithm.
-Reinforcement Learning (RL) operates under the formalization of Markov Decision Process (MDP), $\mathcal{M}=\langle\mathcal{S},\mathcal{A},R,T,\gamma,\rho_{0}\rangle$.
+Our second phase, online fine tuning, further improves the warm-start policy by learning from the data gathered online using the policy trained through BC.
+This phase combines the best aspects of RL and Dataset Aggregation (DAgger) [24] algorithm.
+RL operates under the formalization of Markov Decision Process (MDP), $\mathcal{M}=\langle\mathcal{S},\mathcal{A},R,T,\gamma,\rho_{0}\rangle$.
 $\mathcal{S}$ is the state space and $\mathcal{A}$ denotes the action space.
@@
 $\rho_{0}$ denotes the initial state probability distribution.
-A policy, $\pi:S\rightarrow[0,1]^{|A|}$, is a mapping from states to actions or to a probability distribution over actions.
+A policy, $\pi:S\rightarrow A$ , is a mapping from states to actions or to a probability distribution over actions.
 The objective of RL is to find the policy that optimizes the expected discounted return, $\pi^{*}=\mathbb{E}_{\tau\sim\pi}\left[\sum^{\infty}_{t=0}\gamma^{t}R(s_{t})\right]$.
-DAgger builds upon behavior cloning (BC) by incorporating online interaction with the environment and online querying of the expert.
+DAgger builds upon BC by incorporating online interaction with the environment and online querying of the expert.
 Unlike BC, which trains solely on a fixed dataset of expert demonstrations, DAgger actively collects data from interactions with the environment and solicits expert feedback to augment its training.
 This online improvement process allows DAgger to learn from a more diverse set of experiences, adapt to new situations, and refine the agent’s policy over time, ultimately leading to improved performance in imitation learning tasks.
-In our framework, RL is used to address the sub-optimality and covariance shift problem caused by offline behavior cloning.
-DAgger, on the one hand, is adopted to regulate the trajectory prediction model during reinforcement learning training, ensuring it does not forget the experience learned from behavior cloning.
-Further more, DAgger improves the performance of imitation learning by expanding the diversity of the expert MPC’s demonstrations.
-Given that offline reinforcement learning algorithms require lots of data to effectively cover the entire state space, we choose to fine-tune our pre-trained trajectory prediction model using Proximal Policy Optimization (PPO) [schulman2017proximal] as our RL training algorithm.
-The trajectory prediction model acts as a policy network in the PPO framework.
-DAgger is integrated as a loss term in the RL training, as shown in Equation (3).
-$L_{RL}$ is the standard PPO loss which includes three parts: $L_{policy}$, $L_{entropy}$, and $L_{value}$.
-$L_{imitation}$ is the loss signal from DAgger.
-It represents MSE between the expert MPC’s solution and the warm-started initial guess output from $\pi_{\theta}^{warm}$.
-$\lambda$ is the weight coefficient for the RL loss and imitation loss.
-$L$ $=\lambda\cdot L_{RL}+(1-\lambda)\cdot L_{imitation}$ (3)
-$L_{RL}$ $=\lambda_{1}\cdot L_{policy}+\lambda_{2}\cdot L_{entropy}+\lambda_{3}\cdot L_{value}$
-$xte=distance(pos^{car}_{t},\mathbf{p}_{i}(x_{i},y_{i})^{closest})$ (4)
+In our framework, DAgger and RL address the covariance shift problem caused by offline BC.
+Additionally, RL not only fine-tunes the warm-start solution but also has the unique advantage of directly optimizing MPC computation time.
+We choose to leverage Twin Delayed Deep Deterministic Policy Gradient (TD3) [25] as the RL algorithm to fine-tune our pre-trained trajectory prediction model.
+The trajectory prediction model acts as a policy network in the TD3 framework.
+DAgger is integrated as a term in actor loss in the RL training, as shown in Equation (3).
+$B$ is the batch size and $Q_{\theta}$ is the critic network.
+$L_{imitation}$ is the loss signal from DAgger, representing MSE between the expert MPC’s solution and the warm-started initial guess output from $\pi_{\theta}^{warm}$.
+Note that DAgger is only added when the quality of the initial guess is lower than a certain threshold.
+Because we only need to query the expert MPC when the initial guess is of low quality and calling the expert MPC at each step could be time-consuming.
+$\lambda$ is the weight coefficient between the RL loss and imitation loss.
+$L$ $=\lambda\cdot L_{actor}+(1-\lambda)\cdot L_{imitation}$ (3)
+$L_{\text{actor}}$ $=-\frac{1}{B}\sum_{i=1}^{B}Q_{\theta}(s_{i},\pi^{\text{warm}}(s_{i}))\cdot L_{value}$
 In our driving task, the reward at each step of the RL training is shown in Equation (4).
-The first term is the negative MPC optimization time, and the second term is the negative of the accumulated Cross Track Error ($xte$) over the planning horizon, $H$.
+The first term is the negative MPC optimization time, and the second term is the negative of the Cross Track Error ($xte$) over the planning horizon, $H$.
 The $xte$ is defined in Equation (5).
@@
 This estimate affects both $L_{policy}$, which encourages actions leading to higher rewards, and $L_{value}$, which trains the value function to better estimate cumulative rewards.
-$r_{t}=-time_{MPC}-\sum_{i=t+1}^{t+H}xte(pos_{i}^{car},Traj^{ref})$ (5)
-This reward design helps to optimize the $\pi_{\theta}^{warm}$ by minimizing the MPC running time and by minimizing the difference between the planned trajectory and the reference trajectory, thereby reducing MPC optimization time.
-This helps to address the sub-optimality problem caused by behavior cloning.
+$r_{t}=-time_{MPC}-xte(pos_{t}^{car},Traj^{ref})$ (4)
+$xte=distance(pos^{car}_{t},\mathbf{p}_{i}(x_{i},y_{i})^{closest})$ (5)
+This reward design helps to optimize the $\pi_{\theta}^{warm}$ by minimizing the MPC running time and by maximizing the tracking accuracy.
+This helps to address the sub-optimality problem caused by BC.
 IV EXPERIMENTS
-In this section, we will discuss in detail the setup of the Formula 1 tracks domains, followed by the MPC design in this testing domain.
-Additionally, we will present experimental results from both training and holdout tracks, demonstrating how our algorithm enhances MPC’s performance in terms of both optimization time and tracking accuracy.
+In this section, we will discuss in detail the setup of the Formula 1 track domains, followed by the MPC design in this testing domain.
+Additionally, we will present experimental results from both training and zero-shot tracks, demonstrating how our algorithm enhances MPC’s performance in terms of both optimization time and tracking accuracy.
 IV-A Experiment Setup
-The experiments are done in high-speed Formula 1 tracks [f1tenth_racetracks] which are divided into three training tracks and seven zero-shot testing tracks, as shown in Figure 3(a), Figure 3(b), and Figure 3(c), respectively.
+The experiments are done in high-speed Formula 1 tracks [26] which are divided into three training tracks and seven zero-shot testing tracks, as shown in Figure 3(a), Figure 3(b), and Figure 3(c), respectively.
 The reference trajectory is represented as a set of waypoints in the center of the track.
@@
 The friction between the tire and the road is not considered in our dynamics model, as it depends on multiple factors such as tire pressure, temperature, humidity, and road conditions.
-The path tracking is challenging for traditional gradient-free MPC because the vehicle is running at a high speed and the tracks have multiple sharp turns.
-In this scenario, traditional gradient-free MPC solvers could not optimize the control solution in real time, which makes it unreasonable for real-world application.
 The three tracks used for demonstration collection and training are shown in Figure 3(a).
@@
 Then, the warm-start policy is fine-tuned on the same three tracks using the algorithm introduced in Section III-B.
-As shown in Figure 2, without a specified threshold, optimization continues until reaching the maximum iteration limit.
-In our experiments, we found that the MPC cost reduces subtly during the later optimization iterations.
-As such, it is important to provide the solver with an early-stop criterion.
-We implement the early-stop condition based on a planned trajectory’s accumulated cross-track error ($xte$) being $<0.1m$.
-Additionally, to ensure real-time planning during testing, we cap the maximum optimization iterations at 50, which guarantees that the optimization time is less than 0.08 seconds in worst-case scenarios.
-Details of the MPC design and hyper parameters in MPC and training are presented in Appendix VI-A and VI-B
-(a) Training tracks
-(b) Zero-shot tracks (complex)
-(c) Zero-shot tracks (simple)
-Figure 3: Tracks tested during training and testing.
-IV-B Experiment Results
-During testing, we evaluate on two metrics: the average MPC optimization time (seconds) per step and the average $xte$ (m) per step.
-We benchmark our algorithm’s performance against MPC using all-zero initial guesses and initial guesses derived from the MPC’s solution at the previous step.
-We perform testing on both the training tracks (Figure 3(a)) and the challenging zero-shot tracks (Figure 3(b)).
-The results are shown in Figure 4 and Figure 5, respectively.
-IV-B1 Optimization Time
-On both the training and zero-shot tracks, employing a warm-start policy trained through either offline learning or a combination of offline and online fine-tuning significantly reduces MPC optimization time, falling well within the upper bound of $0.08s$.
-Additionally, the warm-start policy trained with both offline and online fine-tuning achieves better MPC optimization time compared to the policy trained solely through offline behavior cloning, demonstrating the capability of our online fine-tuning algorithm in addressing the suboptimality problem caused by behavior cloning (BC).
-Furthermore, our two-phase training framework achieves more significant improvement in the zero-shot domains, indicating that the online fine-tuning phase alleviates the covariance shift problem caused by BC and enhances the model’s generalizability.
-IV-B2 Control Performance ($xte$)
-In the context of the training tracks, the warm-start policy trained through offline and online fine-tuning outperforms the policy trained solely through offline behavior cloning in terms of tracking accuracy, further illustrating the effectiveness of our online fine-tuning algorithm in addressing the suboptimality problem caused by BC.
-Given that the width of the downscaled track is $2.2m$, achieving an $xte$ of less than $0.3m$ demonstrates precise tracking of the reference path.
-Additionally, our two-phase training framework exhibits more significant improvement in the zero-shot domains, reinforcing the efficacy of the online fine-tuning phase in mitigating the covariance shift problem caused by BC and improving the model’s generalizability.
-Figure 4: Experiment results on training tracks.
-Figure 5: Experiment results on the six zero-shot testing tracks.
-The results also reveal that employing an MPC with either an all-zero initial guess or an initial guess derived from the MPC’s previous solution fails to complete laps on both the training and zero-shot tracks.
-The vehicle consistently deviates from the lane and struggles to navigate sharp turns.
-This limitation arises because the MPC often requires more iterations to optimize the control solution as the vehicle approaches the curves of the track.
-Since the real-time MPC only optimizes the solution for a maximum of 50 iterations at each step, the returned solution lacks the optimization necessary to guide the vehicle through sharp turns effectively.
-This underscores the necessity of a well-informed initial guess to minimize the number of optimization iterations required.
-As the real-time MPC with either an all-zero initial guess or an initial guess derived from the MPC’s previous solution fails to complete laps on both the training and challenging zero-shot tracks, we conducted further testing on four warm start policies on a simple zero-shot track, IMS, as depicted in Figure 3(c).
-The results are summarized in Figure 6.
-Our method demonstrates a significant reduction in MPC optimization time while maintaining high tracking accuracy, showcasing its effectiveness in enhancing MPC performance.
-Figure 6: Experiment results on the simple IMS track.
-As the vehicle route on the track, we further compute the curvature of the track surrounding the vehicle.
-This analysis allows us to illustrate the correlation between the curvature and lateral error ($xte$), depicted in Figure 7.
-It shows that our second phase fine tuning has an advantage on average $xte$ and this advantage has a modest increase when the track is less curvy.
-This outcome underscores the efficacy of our proposed algorithm in enhancing the precision of the planned trajectory.
-Details regarding the method employed to compute the curvature of the track around the vehicle are provided in the Appendix VI-C
-Figure 7: We run the MPC with two types of warm-start policy (BC and ours) on the ”Catalunya” track, where we gather data on the $xte$ and track curvature surrounding the vehicle.
-Subsequently, we depict these data points in the figure above.
-In summary, our empirical results support that:
-  • MPC with initial guesses trained by data-driven methods significantly outperforms those with all-zero or previous solution-derived guesses, crucial for navigating sharp turns.
-  • The warm-start policy trained with our two-phase learning framework reduces MPC optimization time and $xte$ on both training and zero-shot tracks, showing the capability of our online fine-tuning algorithm in addressing the suboptimality problem.
-  • The two-phase training shows superior performance in zero-shot scenarios, indicating better generalizability.
-V LIMITATIONS AND FUTURE WORKS
-While our proposed two-phase learning framework shows promising results in expediting optimization processes and enhancing control performance for robot control tasks, there are several limitations and avenues for future research that merit consideration.
-One limitation of our current experiments is that we have only focused on autonomous driving where the action space is limited to two dimensions (i.e., steering and acceleration).
-Future work could expand into more complex domains, such as robot arm manipulation or quadcopter control, to explore the algorithm’s efficacy in higher-dimensional spaces.
-Additionally, future research could also involve utilizing our proposed algorithm to expedite other iterative control optimization techniques like MPPI.
-Although our algorithm primarily interfaces with MPC, it is versatile enough to serve as a warm start for other iterative control optimization methods like Model Predictive Path Integral (MPPI) control [williams2015model].
-MPPI typically employs stochastic optimization, initializing with a guess of the control sequence distribution, sampling sequences, and iteratively refining the distribution to minimize control cost.
-Our framework can be trained to predict the trajectory Gaussian distribution mean, enhancing MPPI’s performance by offering a superior initial distribution.
-This reduces the iteration count needed for optimization, aligning with early stopping criteria.
-While MPC suits systems with complex dynamics and constraints, we choose it as our focus due to its representativeness in trajectory optimization methods.
-VI CONCLUSIONS
-In this paper, we introduce a novel approach to accelerate MPC optimization by learning a warm-start policy.
-Our two-phase learning framework combines offline behavior cloning and online fine-tuning to provide improved initial guesses for the MPC solver.
-Experimental results on both training and zero-shot tracks demonstrate the effectiveness of our approach in reducing optimization time and enhancing path tracking accuracy.
-Our learning framework integrated with MPC opens up new avenues for improving the efficiency and applicability of trajectory optimization in various dynamic systems.
-APPENDIX
-VI-A MPC Design Details
 MPC is designed to control the acceleration and the steering of the vehicle to track the reference trajectory.
@@
 $xte$ is computed using Equation 5.
-$eth$ denotes the angular disparity between the intended path direction and the current heading of a vehicle in path tracking systems.
-And $eth$ is computed using Equation (8).
+$eth$ is computed using Equation (8), which denotes the angular disparity between the intended path direction and the current heading of a vehicle in path tracking systems.
 $v_{t}^{\text{ref}}$ is the desired speed of the vehicle.
-The last two terms in Equation (7) regulate the rate of change of the steering angle and acceleration to make planned trajectory smoother.
+The last two terms regulate the rate of change of the steering angle and acceleration to make planned trajectory smoother.
 $w_{0},w_{1},w_{2},w_{3},w_{4}$ are the coefficients balancing the importance of each term.
 The planning horizon of the MPC is 25 steps and the planning step $dt$ is 0.02 seconds, which means that the MPC looks 0.5 seconds ahead.
+Since the calculation of $xte$ involves finding the closest point on a reference trajectory to the car’s current position, it introduces discontinuities and non-differentiability into the MPC objective function.
+As a result, only gradient-free, sampling-based solvers are suitable for this problem [27, 28].
+Here we choose to use the gradient-free solver COBYLA [29] in our experiments.
+Furthermore, given the high speed of the vehicle and the presence of multiple sharp turns on the track, gradient-free MPC solvers without a good initial guess of the control solution will struggle to optimize the trajectory in real time, making real-world implementation challenging.
+And we show in Section IV-C gradient-free MPC solver only succeeds in simple IMS track and fails at all of the complex tracks.
 $J_{t}$ $=\sum_{i=t}^{t+H-1}(w_{0}\cdot xte_{i}^{2}+w_{1}\cdot eth_{i}^{2}$ (7)
@@
 $eth=abs\left(yaw^{car}_{t}-\arctan\left({\frac{y_{i+1}-y_{i}}{x_{i+1}-x_{i}}}\right)\right)$ (8)
-VI-B Hyper Parameters in MPC and Training
-Hyperparameters Values
-Weight in MPC cost function $w_{0}$ 2000
-Weight in MPC cost function $w_{1}$ 100
-Weight in MPC cost function $w_{2}$ 60
-Weight in MPC cost function $w_{3}$ 2
-Weight in MPC cost function $w_{4}$ 20
-Weight in MPC cost function $w_{5}$ 300
-Loss weight coefficient $\lambda$ 0.9
-Loss weight coefficient $\lambda_{1}$ 0.5
-Loss weight coefficient $\lambda_{2}$ 0
-Loss weight coefficient $\lambda_{3}$ 0.5
-MPC planning horizon 25
-Length of each time step 0.02 second
-Reference speed $10m/s$
-Wheelbase 2.89 m
-Expert MPC maximum optimization iteration 300
-Real-time MPC maximum optimization iteration 300
-Real time MPC early stop criteria $xte<0.1$
-VI-C Curvature Computation
-Algorithm 3 Curvature Computation
-1:Extract 10 consecutive nearest waypoints ahead of the vehicle, $\mathbf{P}=\{\mathbf{p}_{1},\mathbf{p}_{2},...,\mathbf{p}_{10}\}$. $\mathbf{p}_{i}=(x_{i},y_{i})$ represent the $i$-th waypoint
-2:for $\mathbf{p}_{i-1}$, $\mathbf{p}_{i}$, and $\mathbf{p}_{i+1}$ in $\mathbf{P}$ do
-3:     Compute the vector $\mathbf{v}_{1}=\mathbf{p}_{i}-\mathbf{p}_{i-1}$
-4:     Compute the vector $\mathbf{v}_{2}=\mathbf{p}_{i+1}-\mathbf{p}_{i}$
-5:     Calculate the dot product $\mathbf{v}_{1}\cdot\mathbf{v}_{2}$
-6:     Calculate the magnitudes $\|\mathbf{v}_{1}\|$ and $\|\mathbf{v}_{2}\|$
-7:     Compute the $\cos$ of the angle between the vectors:
-$\cos(\theta_{i})=\left(\frac{\mathbf{v}_{1}\cdot\mathbf{v}_{2}}{\|\mathbf{v}_{1}\|\|\mathbf{v}_{2}\|}\right)$
-8:end for
-9:$Curvature=\sum_{i=2}^{9}\left(1-cos(\theta_{i})\right)$
+We implement the early-stop condition based on a planned trajectory’s accumulated cross-track error ($xte$) being $<0.1m$ to avoid MPC wasting time on optimizing a trajectory that is already good enough.
+Additionally, to ensure real-time planning during testing, we cap the maximum optimization iterations at 50, which guarantees that the optimization time is less than 0.05 seconds in worst-case scenarios.
+As, 20Hz running rate of MPC is validated by [30] for vehicle path tracking.
+(a) Training tracks
+(b) Zero-shot tracks (complex)
+(c) Zero-shot tracks (simple)
+Figure 3: Tracks tested during training and testing.
+IV-B Baselines and Metrics
+During testing, we evaluate on two metrics: the average MPC optimization time (seconds) per step and the average $xte$ (m) per step.
+We compare our algorithm’s performance against five types of initial guess policies: All Zero: all-zero initial guesses, Previous Solution: initial guesses derived from the MPC solution at the previous step, Ours w/o BC: initial guesses from a warm-start policy trained exclusively with an online training algorithm, without offline training, Ours w/o RL: initial guesses from a warm-start policy trained solely with offline BC, and Ours w/o DAgger: initial guesses from a warm-start policy trained with our two-phase learning algorithm but without the use of DAgger.
+To ensure a fair comparison between the Ours w/o BC and warm-start policy trained by our two-phase learning algorithm, we extend the training time of Ours w/o BC by an additional hour to account for the data collection time in the offline BC phase.
+IV-C Experiment Results
+We perform testing on both the training tracks (Figure 3(a)) and the challenging zero-shot tracks (Figure 3(b)).
+The results are shown in Figure 4 and Figure 5, respectively.
+On both the training and zero-shot tracks, employing a warm-start policy trained through either offline learning or a combination of offline and online fine-tuning significantly reduces MPC optimization time and improves tracking accuracy.
+Besides, the warm-start policy trained with both offline and online fine-tuning achieves better MPC optimization time and tracking accuracy compared to the policy trained solely through offline BC, demonstrating the capability of our online fine-tuning algorithm in addressing the suboptimality covariance shift problem caused by BC.
 […]
+The results also reveal that initialize an MPC with either All Zeros or Previous Solution fails to complete laps on both the training and zero-shot tracks.
+The vehicle consistently deviates from the lane when facing any sharp turns.
+This limitation arises because the MPC often requires more iterations to optimize the control solution as the vehicle approaches the curves of the track.
+Since the real-time MPC only optimizes the solution for a maximum of 50 iterations at each step, the returned solution lacks the optimization necessary to guide the vehicle through sharp turns effectively.
+This underscores the necessity of a well-informed initial guess to minimize the number of optimization iterations required.
+Figure 4: Experiment results on training tracks. The asterisk (*) implies statistical significance: * $p<0.05$, ** $p<0.01$, *** $p<0.001$, n.s. means no significant.
+Figure 5: Experiment results on the six zero-shot testing tracks. The asterisk (*) implies statistical significance: * $p<0.05$, ** $p<0.01$, *** $p<0.001$, n.s. means no significant.
+Furthermore, the Ours w/o RL algorithm also fails on both training and zero-shot domains.
+This is because the absence of offline BC requires the algorithm to explore the solution space from scratch, causing the online learning algorithm to take significantly longer to converge to a satisfactory solution.
+Given the same amount of training time, the quality of the policy learned by Ours w/o RL lags behind that of the policy learned through the two-phase training algorithm.
+This demonstrates the effectiveness of offline BC in accelerating online training.
+To further inspect the relative performance of our approach when our baselines performed better, i.e.
+on simpler tracks, we conducted further testing on six warm-start policies on a simpler zero-shot track, IMS, as depicted in Figure 3(c),   similar to the path-tracking scenario in [31].
+The results are summarized in Figure 6.
+Our method demonstrates a significant reduction in MPC optimization time while maintaining high tracking accuracy, showcasing its effectiveness in enhancing MPC performance.
+Figure 6: Experiment results on the simple IMS track.
+In summary, our empirical results support that:
+  • MPC with warm-start policy trained by either offline learning or a combination of offline and online fine-tuning significantly outperforms those with all-zero or previous solution-derived guesses, crucial for navigating sharp turns.
+  • The offline training algorithm significantly warm-starts the online training algorithm and make it take much less training time to get a satisfying performance.
+  • The warm-start policy trained with our two-phase learning framework reduces MPC optimization time and $xte$ on both training and zero-shot tracks, showing the capability of our online fine-tuning algorithm in addressing the suboptimality problem.
+  • The two-phase training algorithm outperforms in zero-shot scenarios, demonstrating superior generalizability.
+V LIMITATIONS AND FUTURE WORKS
+While our proposed two-phase learning framework shows promising results in expediting optimization processes and enhancing control performance for robot control tasks, there are still limitations and avenues for future research that merit consideration.
+Future work could expand into more complex domains, such as robot arm manipulation or quadcopter control, to explore the algorithm’s efficacy in higher-dimensional spaces.
+Additionally, future research could also involve utilizing our proposed algorithm to expedite other iterative control optimization techniques like MPPI.
+Although our algorithm primarily interfaces with MPC, it is versatile enough to serve as a warm start for other iterative control optimization methods like Model Predictive Path Integral (MPPI) control [32].
+MPPI typically employs stochastic optimization, initializing with a guess of the control sequence distribution, sampling sequences, and iteratively refining the distribution to minimize control cost.
+Our framework can be trained to predict the trajectory Gaussian distribution mean, enhancing MPPI’s performance by offering a superior initial distribution.
+This reduces the iteration count needed for optimization, aligning with early stopping criteria.
+While MPC suits systems with complex dynamics and constraints, we choose it as our focus due to its representativeness in trajectory optimization methods.
+VI CONCLUSIONS
+In this paper, we introduce a novel approach to accelerate MPC optimization by learning a warm-start policy.
+Our two-phase framework combines offline BC and online fine-tuning to provide better initial guesses for the MPC solver.
+Experiments on training and zero-shot tracks demonstrate the effectiveness of our approach in reducing optimization time and improving path tracking accuracy.
+This integration of learning with MPC enhances the efficiency and applicability of trajectory optimization in dynamic systems.

### v3

@@
 Abstract
-Optimization for robot control tasks, spanning various methodologies, includes Model Predictive Control (MPC).
-However, the complexity of the system, such as non-convex and non-differentiable cost functions and prolonged planning horizons often drastically increases the computation time, limiting MPC’s real-world applicability.
-Prior works in speeding up the optimization have limitations on optimizing MPC running time directly and generalizing to hold out domains.
-To overcome this challenge, we develop a novel framework aiming at expediting optimization processes directly.
-In our framework, we combine offline self-supervised learning and online fine-tuning to improve the control performance and reduce optimization time.
-We demonstrate the success of our method on a novel and challenging Formula 1 track driving task.
-Comparing to single-phase training, our approach achieves a $19.4\%$ reduction in optimization time and a $6.3\%$ improvement in tracking accuracy on zero-shot tracks.
+Model Predictive Control (MPC) is widely used in robot control by optimizing a sequence of control outputs over a finite-horizon.
+Computational approaches for MPC include deterministic methods (e.g., iLQR and COBYLA), as well as sampling-based methods (e.g., MPPI and CEM).
+However, complex system dynamics and non-convex or non-differentiable cost terms often lead to prohibitive optimization times that limit real-world deployment.
+Prior efforts to accelerate MPC have limitations on: (i) reusing previous solutions fails under sharp state changes and (ii) pure imitation learning does not target compute efficiency directly and suffers from suboptimality in the training data.
+To address these, We propose a warm-start framework that learns a policy to generate high-quality initial guesses for MPC solver.
+The policy is first trained via behavior cloning from expert MPC rollouts and then fine-tuned online with reinforcement learning to directly minimize MPC optimization time.
+We empirically validate that our approach improves both deterministic and sampling-based MPC methods, achieving up to 21.6% faster optimization and 34.1% more tracking accuracy for deterministic MPC in Formula 1 track path-tracking domain, and improving safety by 100%, path efficiency by 12.8%, and steering smoothness by 7.2% for sampling-based MPC in obstacle-rich navigation domain.
+These results demonstrate that our framework not only accelerates MPC but also improves overall control performance.
+Furthermore, it can be applied to a broader range of control algorithms that benefit from good initial guesses.
 I INTRODUCTION
 Algorithms that optimize control outputs iteratively like MPC have been widely adopted to control dynamic systems, such as autonomous vehicles [1, 2, 3, 4], aircraft [5, 6], humanoid robots [7], etc.
-However, MPC in real-world applications is challenging due to computational time bottlenecks [8, 9, 10, 11], particularly with complex systems featuring non-convex and non-differentiable cost functions and prolonged planning horizons.
-[12, 13].
-Addressing this challenge is paramount for enhancing the practical utility of MPC in various domains.
+While gradient-based MPC is highly effective when the dynamics and cost functions are smooth and differentiable, applying MPC in real-world settings remains challenging.
+Computational bottlenecks often arise in systems with non-convex or non-differentiable costs and long planning horizons, where gradient-free solvers are typically employed [8, 9, 10, 11].
 Figure 1: Overview of our proposed algorithm.
@@
 During testing, the proposed framework is evaluated on both training tracks and challenging zero-shot tracks, as demonstrated in the third block.
-Warm starting in robot control optimization involves initializing the solver with a closer-to-optimal solution, speeding up the iterative optimization process.
-The intuition behind warm starting is that, at each iteration, the solver refines its solution based on previous iterations.
-As shown in Figure 2, a closer initial guess reduces the distance the optimization algorithm needs to traverse to converge to the optimal solution.
-By starting closer to the optimum, the solver can often avoid lengthy exploration of the solution space, leading to faster convergence and reduced computational effort.
+Dynamic replanning in MPC requires fast computation to operate in real time, but standard MPC without warm-starting often fails to achieve sufficient speed in challenging domains [12, 13].
+Warm starting addresses this by initializing the solver with a solution close to the optimum, reducing the search space and accelerating convergence.
 In this paper, we choose to test our proposed algorithm to warm start Model Predictive Control (MPC), which optimizes a trajectory over a predefined horizon.
-To provide MPC with a high quality initial guess, various approaches have been adopted.
+Prior works have employed heuristic strategies to warm-start MPC, but these methods often struggle with scalability and fail to adapt effectively to sudden changes in system states.
 Traditionally, one common technique involves utilizing the MPC solution from the previous sampling instance as the initial guess for the current control step [12, 14].
 However, this approach falls short when faced with sudden state changes (e.g., the vehicle approaches a sharp turn).
-Another method involves maintaining a memory buffer to store historical MPC solutions, from which a suitable initial guess can be retrieved for future planning steps [10, 16].
-However, this approach may lack zero-shot generalizability.
-As the system’s state becomes heterogeneous, the size of the memory buffer needs to increase, thus causing time-consuming operations to searching for the initial guess within the buffer.
-Alternatively Klaučo et al.
-propose using a k-NN classifier to classify the solution space into active sets for the solver to search for a solution [17].
-However, this method lacks precise warm-start guesses and struggles with complex control tasks involving long planning horizons and heterogeneous observations.
-Moreover, it’s limited to strictly convex quadratic programs, reducing its applicability to real-world problems with non-linear or non-differentiable objectives and dynamics.
-More recent approaches have adopted supervised learning to boost the performance of model predictive optimization algorithms.
-[18, 19] adopts behavior cloning (BC) to learn the distribution of the initial guesses for the solver.
-However, supervised learning cannot directly learn to minimize the computation time of MPC.
+Another method involves maintaining a memory buffer to store and leverage historical MPC solution [10, 15].
+However, this approach fails to generalize to unseen states and scales.
+Traditional machine learning approach has also been adopted to learn a warm-start for MPC, however, it suffers from inaccurate initialization.
+Klaučo et al.
+propose using a k-NN classifier to classify the solution space into active sets for the solver to search for a solution [16], but this only produces coarse guesses.
+These shortcomings highlight the need for a general warm-starting framework that directly learns to generate high-quality initial guesses, scales to complex domains, and improves solver efficiency under sharp dynamics and diverse conditions.
+More recent approaches have attempted to predict effective warm-starts for model predictive optimization by either behavior cloning (BC) historical trajectories [17, 18] or training reinforcement learning (RL) policies for trajectory generation [19].
+While BC is fast and straightforward to train, it cannot directly learn to minimize the optimization time of MPC.
 Instead, all supervised learning can do is to imitate the output of the solver.
-Besides, BC also suffers from covariance shift, which degrades model’s performance on zero-shot scenarios [20].
-[21] used reinforcement learning (RL) to train a warm-start policy to help quadrotor track simple zig-zag trajectories.
-Yet, RL algorithms are too data hungry to learn an optimal policy from scratch, posing challenge to complex, long-horizon tasks [22].
-Furthermore, [21] is still not directly learning to minimize the computation time of MPC.
-Thus, there is a pressing need for innovative approaches that can effectively address the challenges posed by the inherent complexity and non-linearity of real-world control systems.
-In this paper, we propose a two-phase learning framework to learn a warm-start policy that provides a better initial guess for iterative control optimization algorithms (such as MPC) to reduce the optimization time.
-Our proposed algorithm is depicted in Figure 1, which includes a two-phase training framework and an additional zero-shot testing phase.
-The two training phase works reciprocally to each other.
-On the one hand, the online-fine tuning phase uses RL to address the suboptimality and covariance shift problem of the offline BC by directly optimize the computation time of MPC.
-On the other hand, the offline BC provides online learning algorithm with a good starting point to expedite the training process.
-Moreover, another advantage of using a learned policy to initialize the MPC, rather than for end-to-end control, is that it preserves the original definition of the control problem.
-The learned policy only affects the starting point, while MPC ensures the control solution satisfies system constraints, leading to safer operation.
+Worse yet, BC suffers from covariance shift, leading to poor generalization in zero-shot scenarios [20].
+In contrast, RL is capable of directly optimizing performance objectives but requires extensive interaction data and struggles to learn from scratch [21].
+As such, there remains a need to bridge these two paradigms—combining the efficiency of BC with the adaptability of RL—to effectively warm-start MPC across diverse solver families.
+To overcome the limitations of prior work, we propose a learning framework that trains a policy to generate high-quality initial guesses for optimization-based control algorithms (e.g., MPC).
+The framework combines offline behavior cloning (BC) with online reinforcement learning (RL) fine-tuning and an additional zero-shot testing phase, as depicted in Fig.
+1.
+The two training phases work synergistically.
+The online-fine tuning phase uses RL to address the suboptimality and covariance shift problem of the offline BC by directly optimizing the optimization time of MPC.
+At the same time, the offline BC provides the online learning algorithm with a good starting point to expedite the training process.
+Moreover, initializing MPC with a learned policy rather than replacing MPC with an end-to-end controller preserves the original control formulation.
+The policy only sets the starting point, while MPC enforces system constraints to ensure safer operation.
+We demonstrate the generality of our approach by evaluating it across both deterministic MPC solver (COBYLA [22]) and sampling-based MPC solver (MPPI [23]).
 Our key contributions are:
   1. 1.
-    Propose a two-phase learning framework featuring offline training and online fine-tuning to learn a warm-start policy that provides the iterative control optimization algorithm with higher-quality initial guesses, aiding the command of a high-speed vehicle on multiple novel, challenging Formula 1 tracks in real-time where traditional warm start methods all failed.
+    We propose a novel two-phase learning framework to warm-start optimization-based control algorithms (e.g., MPC), enhancing both efficiency and robustness.
   2. 2.
-    Empirically evaluate our proposed two-phase learning framework and show the online fine-tuning phase helps the iterative control optimization algorithm achieve a $19.4\%$ higher performance in optimization time and a $6.3\%$ higher performance in tracking accuracy on challenging zero-shot tracks comparing to single-phase training.
-Figure 2: An illustrative demo showing how a better initial guess improves MPC optimization time.
-II Preliminaries
-II-A Model Predictive Control
-MPC is formulated as an optimization task, where the objective is to minimize a cost function while satisfying constraints on the system’s inputs and outputs over a finite horizon, H.
-This optimization problem is typically solved iteratively at each time step, generating a sequence of control inputs that steer the system towards a desired state while considering future predictions of its behavior.
-When MPC observes a new state, $x_{t}$, at time step, $t$, it optimizes a sequence of control inputs, $U$, over the planning horizon by iteratively minimizing a cost function , $J$ while respecting system dynamics and constraints.
-Then, only the first action in the action sequence is applied to the controlled system, which leads the system to the next state.
-The optimization process is performed again at the next state.
-The MPC control law is formulated as shown below in Equations (1):
-[…]
-subject $\text{subject to}\quad x_{t+i+1}=f(x_{t+i},u_{t+i})$ (1b)
+    We empirically evaluate the framework on high-speed vehicle control tasks ( deterministic MPC), achieving $21.6\%$ faster optimization and $34.1\%$ higher tracking accuracy on zero-shot Formula 1 tracks compared to single-phase training.
+  3. 3.
+    We further test the framework on path-planning tasks (sampling-based MPC), showing $100\%$ improved safety, $12.8\%$ higher path efficiency, and $7.2\%$ smoother steering on zero-shot tracks compared to single-phase training.
+II Preliminaries - Model Predictive Control
+MPC is formulated as a finite-horizon optimization problem that minimizes a cost function while satisfying system dynamics and constraints.
+The MPC control law is formulated in Eqs.(1).
+At each time step $t$, MPC observes the current state $x_{t}$ and computes an optimal sequence of control inputs $U=[u_{t},\ldots,u_{t+H-1}]$ over a planning horizon $H$ by minimizing a cost function $J$ (Eq.
+(1a)).
+The system dynamics $x_{t+i+1}=f(x_{t+i},u_{t+i})$ (Eq.
+(1b)) predict how control inputs influence future states.
+Eq.
+(1c) and Eq.
+(1d) specify the constraints for the control inputs and states.
+After solving the optimization, only the first control input $u_{t}$ is applied, and the process repeats at the next time step using the updated state.
+$J=\underset{U}{\text{minimize}}\sum_{i=0}^{H-1}l(x_{t+i},u_{t+i})$ (1a)
+$\text{subject to}\quad x_{t+i+1}=f(x_{t+i},u_{t+i})$ (1b)
 $U=[u_{t},...,u_{t+H-1}]\in\mathcal{U}_{j}\ \text{for all}\ j=1,\dots,n_{c_{u}}$ (1c)
 $X=[x_{t},...,x_{t+H}]\in\mathcal{X}_{j}\ \text{for all}\ j=1,\dots,n_{c_{x}}$ (1d)
-Equation (1a) formulates the objective function of the MPC.
-$H$ denotes the planning horizon, $t$ is the current step, and, $x_{t+i}$ is the system state at the $t+i$ step.
-Given the state, $x_{t+i}$ and the control input, $u_{t+i}$, the system dynamics model predicts the next state, $x_{t+i+1}$, using the dynamics model (1b).
-Equation (1c) and Equation (1d) are the constraints for the control inputs and states.
-$n_{c_{u}}$ and $n_{c_{x}}$ are the number of constraints on control inputs and states respectively.
 III Method
-In this section, we discuss in detail the proposed two-phase training framework.
+In this section, we discuss in detail the proposed two-phase training framework to learn a warm-start policy that reduces MPC solver runtime while maintaining tracking performance..
 In the first phase, we run the MPC to collect the expert demonstrations, which are represented as state-action pairs.
 Then, we use BC to train a warm-start policy to mimic the expert MPC’s solution, as shown in Algorithm 1.
-The output of warm-start policy is utilized as an initial guess to warm start the MPC and reduce the optimization time.
+The output of warm-start policy is utilized as an initial guess to warm start the MPC.
 In the second phase, we load the pre-trained trajectory prediction model into an online training framework and fine tune the warm-start policy to address the suboptimality problem caused by BC and improve the model’s generalizability.
@@
 Algorithm 1 Offline Training
-1:Input: expert MPC $\pi^{\text{MPC}}$ with planning horizon $H$ and maximum optimization iteration $N_{expert}$ and all-zero vector $\vec{0}$ as initial guess, environment transition $T$, number of state-action pairs to collect $N$
+1:Input: expert MPC $\pi^{\text{MPC}}$ with planning horizon $H$ and maximum optimization iteration $N_{expert}$ and all-zero vector $\vec{0}$ as initial guess, environment transition $T$, number of state-action pairs to collect NN
 2:Initialize neural network policy $\pi_{\theta}^{\text{warm}}$
-3:$t\leftarrow 0,s\leftarrow s_{0},\mathcal{D}=\emptyset$
+[…]
 4:while $t<N$ do
@@
 9:end while
-10:Train $\pi_{\theta}^{warm}$ with Equation (2) and $\mathcal{D}$
+10:Train $\pi_{\theta}^{warm}$ with Eq. (2) and $\mathcal{D}$
+III-A Offline Training
+At this phase, we implement an expert MPC to collect a dataset $\mathcal{D}$ containing $N$ state-action pairs.
+The expert MPC, $\pi^{MPC}$, controls the agent to complete the task without a warm-started initial guess.
+At each step, $\pi^{MPC}$ observes a state $s$, takes an all-zero vector as the initial guess, and optimizes it to output $(u_{t},u_{t+1},\cdots,u_{t+H-1})$ (line 5).
+The state-action pair $(s,u_{t},u_{t+1},\cdots,u_{t+H-1})$ is stored in $\mathcal{D}$ (line 6).
+The first action is then applied, transitioning the system to the next state based on environment transition $T$ (line 7–8).
+During data collection, $\pi^{MPC}$ runs enough iterations to ensure high control performance by disabling the “early stop.”
+We design our warm-start policy, $\pi_{\theta}^{warm}$, as a multi-layer perceptron with ReLU activation function [24].
+Given the current vehicle state, the $\pi_{\theta}^{warm}$ predicts a sequence of actions that serves as the initial guess of the MPC to warm start the optimization process.
+In the offline training phase, we utilize BC to train $\pi_{\theta}^{warm}$ (line 10), where BC learns a control policy $\pi_{\theta}$ from pre-collected MPC demonstrations by minimizing the mean squared error (MSE) between predicted and demonstrated actions, as shown in Eq.
+(2).
+$\theta^{*}=\min_{\theta}{\sum_{\tau^{i}\in\mathcal{D}}\sum_{t=0}^{T}{(\pi_{\theta}^{warm}(s^{i})_{t}-u_{t}^{i})^{2}}}$ (2)
 Algorithm 2 Online Fine Tuning
@@
 5:  $pos_{i+1}^{car}=M_{dynamics}((s_{i},\hat{u}_{i}))|_{i=t}^{H-1}$
-6:     Compute accumulated $xte$ using Equation (5)
+6:  Evaluate the quality of the planned trajectory
 7:  $U_{t}^{\text{MPC}}=(u_{t},u_{t+1},...,u_{t+H-1})\leftarrow\pi^{\text{MPC}}_{\text{fast}}(s,\hat{U}_{t})$
-8:     Calculate reward in RL using Equation (4)
+8:  Calculate reward in RL using Eq. (8)
 9:  $L_{imitation}=MSE(U_{t}^{MPC},U_{t}^{guess})$
-10:     Compute training loss $L$ in Equation (3)
+10:  Compute training loss $L$ in Eq. (3)
 11:  $s\leftarrow T(s,u_{t})$
@@
 13:end for
-III-A Offline Training
-At this phase, we implement an expert MPC to collect a dataset $\mathcal{D}$ containing $N$ state-action pairs first.
-The expert MPC, $\pi^{MPC}$ aims to control the agent to finish the task without relying on a warm-started initial guess for the control action sequence.
-At each step, $\pi^{MPC}$ observes a state $s$ and takes an all-zero vector as the initial guess and optimizes that initial guess to output $(u_{t},u_{t+1},\cdots,u_{t+H-1})$ (line 5).
-Then the state-action pair, $(s,u_{t},u_{t+1},\cdots,u_{t+H-1})$ is stored in the dataset $\mathcal{D}$ (line 6).
-The first action in the action sequence is applied to the system, which leads the system to the next state based on the environment transition $T$ (line 7-8).
-During the data collection stage, $\pi^{MPC}$ optimizes the initial guess for enough iterations to ensure it achieves a high control performance by disabling the “early stop” for the expert MPC.
-We design our warm-start policy, $\pi_{\theta}^{warm}$, as a multi-layer perceptron with ReLU activation function [23].
-Given the current state of the vehicle, the $\pi_{\theta}^{warm}$ predicts a sequence of actions which then serves as the initial guess of the MPC to warm start the optimization process.
-In the offline training phase, we utilize BC to train $\pi_{\theta}^{warm}$ (line 10).
-BC is a simple yet effective algorithm to learn from demonstrations.
-The demonstration is a set of trajectories: $\mathcal{D}=\{\tau_{i}\}$.
-BC learns a control policy, $\pi_{\theta}$, by minimizing the Mean Squared Error (MSE) in the demonstration set, shown in Equation (2).
-Our algorithm utilizes pre-collected MPC data to perform BC as a warm start for the MPC to improves the MPC’s optimization time.
-$\theta^{*}=\min_{\theta}{\sum_{\tau^{i}\in\mathcal{D}}\sum_{t=0}^{T}{(\pi_{\theta}^{warm}(s^{i})_{t}-u_{t}^{i})^{2}}}$ (2)
 III-B Online Fine Tuning
-Our second phase, online fine tuning, further improves the warm-start policy by learning from the data gathered online using the policy trained through BC.
-This phase combines the best aspects of RL and Dataset Aggregation (DAgger) [24] algorithm.
+Our second phase, online fine-tuning, addresses the shortcomings of BC by combining the strengths of RL and Dataset Aggregation (DAgger) [25].
+RL enables us to directly optimize the MPC computation time.
+Meanwhile, DAgger mitigates the distribution shift problem by augmenting the training set with additional expert demonstrations.
+DAgger builds upon BC by incorporating online interaction with the environment and online querying of the expert.
+Unlike BC, which trains solely on a fixed dataset of expert demonstrations, DAgger actively collects data from interactions with the environment and solicits expert feedback to augment its training.
+This online improvement process allows DAgger to learn from a more diverse set of experiences, adapt to new situations, and refine the agent’s policy over time, ultimately leading to improved performance in imitation learning tasks.
 RL operates under the formalization of Markov Decision Process (MDP), $\mathcal{M}=\langle\mathcal{S},\mathcal{A},R,T,\gamma,\rho_{0}\rangle$.
@@
 The objective of RL is to find the policy that optimizes the expected discounted return, $\pi^{*}=\mathbb{E}_{\tau\sim\pi}\left[\sum^{\infty}_{t=0}\gamma^{t}R(s_{t})\right]$.
-DAgger builds upon BC by incorporating online interaction with the environment and online querying of the expert.
-Unlike BC, which trains solely on a fixed dataset of expert demonstrations, DAgger actively collects data from interactions with the environment and solicits expert feedback to augment its training.
-This online improvement process allows DAgger to learn from a more diverse set of experiences, adapt to new situations, and refine the agent’s policy over time, ultimately leading to improved performance in imitation learning tasks.
-In our framework, DAgger and RL address the covariance shift problem caused by offline BC.
-Additionally, RL not only fine-tunes the warm-start solution but also has the unique advantage of directly optimizing MPC computation time.
-We choose to leverage Twin Delayed Deep Deterministic Policy Gradient (TD3) [25] as the RL algorithm to fine-tune our pre-trained trajectory prediction model.
-The trajectory prediction model acts as a policy network in the TD3 framework.
-DAgger is integrated as a term in actor loss in the RL training, as shown in Equation (3).
+DAgger is integrated as a term in actor loss in the RL training, as shown in Eq.
+(3).
 $B$ is the batch size and $Q_{\theta}$ is the critic network.
 $L_{imitation}$ is the loss signal from DAgger, representing MSE between the expert MPC’s solution and the warm-started initial guess output from $\pi_{\theta}^{warm}$.
-Note that DAgger is only added when the quality of the initial guess is lower than a certain threshold.
-Because we only need to query the expert MPC when the initial guess is of low quality and calling the expert MPC at each step could be time-consuming.
+DAgger loss is only added when the quality of the initial guess is lower than a certain threshold.
+To reduce the overhead of repeatedly querying the expert MPC, $L_{imitation}$ is only applied when the initial guess falls below a certain threshold.
+This accelerates RL training while maintaining regularization from expert demonstrations.
 $\lambda$ is the weight coefficient between the RL loss and imitation loss.
@@
 $L_{\text{actor}}$ $=-\frac{1}{B}\sum_{i=1}^{B}Q_{\theta}(s_{i},\pi^{\text{warm}}(s_{i}))\cdot L_{value}$
-In our driving task, the reward at each step of the RL training is shown in Equation (4).
-The first term is the negative MPC optimization time, and the second term is the negative of the Cross Track Error ($xte$) over the planning horizon, $H$.
-The $xte$ is defined in Equation (5).
-At each step, $\pi_{\theta}^{warm}$ will output an action sequence of length $H$ based on the current state vector, $s_{t}$.
-Then, the dynamics model, $M_{dynamics}$, is used to calculate the future positions of the agent.
-The $xte$ is computed between the reference trajectory and the future positions of the vehicle, $pos_{i}^{car}$.
-The reward obtained at each step directly influences $L_{RL}$ by contributing to the calculation of the advantage estimate, which measures the discrepancy between the observed reward and the expected value of the state-action pair.
-This estimate affects both $L_{policy}$, which encourages actions leading to higher rewards, and $L_{value}$, which trains the value function to better estimate cumulative rewards.
-$r_{t}=-time_{MPC}-xte(pos_{t}^{car},Traj^{ref})$ (4)
-$xte=distance(pos^{car}_{t},\mathbf{p}_{i}(x_{i},y_{i})^{closest})$ (5)
-This reward design helps to optimize the $\pi_{\theta}^{warm}$ by minimizing the MPC running time and by maximizing the tracking accuracy.
-This helps to address the sub-optimality problem caused by BC.
-IV EXPERIMENTS
-In this section, we will discuss in detail the setup of the Formula 1 track domains, followed by the MPC design in this testing domain.
-Additionally, we will present experimental results from both training and zero-shot tracks, demonstrating how our algorithm enhances MPC’s performance in terms of both optimization time and tracking accuracy.
-IV-A Experiment Setup
-The experiments are done in high-speed Formula 1 tracks [26] which are divided into three training tracks and seven zero-shot testing tracks, as shown in Figure 3(a), Figure 3(b), and Figure 3(c), respectively.
-The reference trajectory is represented as a set of waypoints in the center of the track.
-The track size is downscaled to a 10:1 ratio to make each lap a reasonable length.
-The reference speed for the vehicle in the expert MPC is $10m/s$.
-Given the downscaled size of the track data, the vehicle could actually speed up to $100m/s$ in the 1:1 scale track.
-The friction between the tire and the road is not considered in our dynamics model, as it depends on multiple factors such as tire pressure, temperature, humidity, and road conditions.
-The three tracks used for demonstration collection and training are shown in Figure 3(a).
-The six zero-shot tracks are not presented to the vehicle before testing to demonstrate our algorithm’s zero-shot generalizability.
-During training, the expert MPC is first rolled out on training tracks to collect demonstrations for offline training introduced in Section III-A.
-Then, the warm-start policy is fine-tuned on the same three tracks using the algorithm introduced in Section III-B.
-MPC is designed to control the acceleration and the steering of the vehicle to track the reference trajectory.
-The dynamics model of the vehicle, $M_{dynamics}$, is shown in Equation (6).
-$(x_{t}^{car},y_{t}^{car})$ is the global position of the vehicle at time step $t$.
-$v_{t}$ and $yaw_{t}$ are the speed ($m/s$) and yaw angle ($rad$) of the vehicle at time step $t$.
-$a_{t}$ and $\theta_{t}^{\text{steering}}$ are acceleration ($m/s^{2}$) and steering angle ($rad$) inputs to the vehicle at time step $t$.
-$L$ is the wheelbase of the vehicle.
-Here we use Ford Mustang’s dimensions, $L=2.89m$.
-And $dt=0.02$ is the length of each time step which represents the duration between successive updates of control inputs and system states.
-$x_{t+1}^{car}=x_{t}^{car}+v_{t}\cdot\cos(yaw_{t})\cdot dt$ (6a)
-$y_{t+1}^{car}=y_{t}^{car}+v_{t}\cdot\sin(yaw_{t})\cdot dt$ (6b)
-$yaw_{t+1}=yaw_{t}+\frac{v_{t}}{L}\cdot\tan(\theta_{t}^{\text{steering}})\cdot dt$ (6c)
-$v_{t+1}=v_{t}+a_{t}\cdot dt$ (6d)
-The MPC objective function is composed of five parts as shown in Equation (7).
-The first two terms are $xte$ and Error in Heading ($eth$).
-$xte$ is computed using Equation 5.
-$eth$ is computed using Equation (8), which denotes the angular disparity between the intended path direction and the current heading of a vehicle in path tracking systems.
-$v_{t}^{\text{ref}}$ is the desired speed of the vehicle.
+IV EXPERIMENTS SETUP
+In this section, we discuss in detail the experimental setup for both deterministic and sampling-based MPC.
+For deterministic MPC, we evaluate the path-tracking problem on high-speed Formula 1 tracks [26], where minimizing optimization time is critical for maintaining performance.
+For sampling-based MPC, we choose the navigation problem in environments without a pre-defined reference path, requiring the vehicle to generate its own collision-free trajectory.
+IV-A Deterministic MPC
+IV-A1 MPC Details
+The Formula 1 domain consists of three training and seven zero-shot testing tracks, as shown in Fig.
+2(a) and Fig.
+2(b).
+The training tracks are used for demonstration collection and policy learning, while the zero-shot tracks remain unseen during training to test generalization.
+Each track’s reference trajectory is defined by waypoints along the centerline and scaled down 10:1 to make each lap a reasonable length.
+The friction between the tire and the road is not considered in our dynamics model.
+(a) Training tracks
+(b) Zero-shot tracks
+Figure 2: Training and Testing maps used in the Formula 1 track path-tracking domain.
+$x_{t+1}^{car}=x_{t}^{car}+v_{t}\cdot\cos(yaw_{t})\cdot dt+\epsilon_{1}$ (4a)
+$y_{t+1}^{car}=y_{t}^{car}+v_{t}\cdot\sin(yaw_{t})\cdot dt+\epsilon_{2}$ (4b)
+$yaw_{t+1}=yaw_{t}+\frac{v_{t}}{L}\cdot\tan(\theta_{t}^{\text{steering}})\cdot dt+\epsilon_{3}$ (4c)
+$v_{t+1}=\begin{cases}v_{t}+a_{t}\,\Delta t+\epsilon_{4},&\text{if }v_{t}+a_{t}\,\Delta t<10,\\ 10,&\text{otherwise.}\end{cases}$ (4d)
+$\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\sigma^{2}I),\quad\boldsymbol{\epsilon}=[\epsilon_{1},\epsilon_{2},\epsilon_{3},\epsilon_{4}]^{\top}$ (4e)
+MPC controls vehicle acceleration and steering to follow the reference trajectory using the dynamics model $M_{dynamics}$ in Eq.
+(4).
+$(x_{t}^{car},y_{t}^{car})$ denote the vehicle’s global position; $v_{t}$ and $yaw_{t}$ are its speed ($m/s$) and yaw angle (radrad); $a_{t}$ and $\theta_{t}^{\text{steering}}$ represent acceleration ($m/s^{2}$) and steering angle (radrad).
+The wheelbase is L=2.89mL=2.89~m, and the time step is dt=0.02sdt=0.02~s.
+Gaussian noise $\epsilon_{1}$ - $\epsilon_{4}$ with $\sigma=0.01$ is added to model uncertainty.
+The MPC objective function is composed of five parts as shown in Eq.
+(5).
+The first two terms are Cross Track Error (xtexte) and Error in Heading (etheth) computed using Eq.
+(6).
+and Eq.
+(7) respectively.
+etheth denotes the angular disparity between the intended path direction and the current heading of a vehicle in path tracking systems.
+$v_{t}^{\text{ref}}=10m/s$ is the desired speed of the vehicle.
 The last two terms regulate the rate of change of the steering angle and acceleration to make planned trajectory smoother.
 $w_{0},w_{1},w_{2},w_{3},w_{4}$ are the coefficients balancing the importance of each term.
-The planning horizon of the MPC is 25 steps and the planning step $dt$ is 0.02 seconds, which means that the MPC looks 0.5 seconds ahead.
-Since the calculation of $xte$ involves finding the closest point on a reference trajectory to the car’s current position, it introduces discontinuities and non-differentiability into the MPC objective function.
-As a result, only gradient-free, sampling-based solvers are suitable for this problem [27, 28].
-Here we choose to use the gradient-free solver COBYLA [29] in our experiments.
-Furthermore, given the high speed of the vehicle and the presence of multiple sharp turns on the track, gradient-free MPC solvers without a good initial guess of the control solution will struggle to optimize the trajectory in real time, making real-world implementation challenging.
-And we show in Section IV-C gradient-free MPC solver only succeeds in simple IMS track and fails at all of the complex tracks.
+The planning horizon of the MPC is 25 steps and the planning step $dt$ is 0.02 seconds.
 $J_{t}$ $=\sum_{i=t}^{t+H-1}(w_{0}\cdot xte_{i}^{2}+w_{1}\cdot eth_{i}^{2}$ (5)
@@
 $\quad+w_{4}\cdot(throttle_{i}-throttle_{i-1})^{2})$
+$xte=distance(pos^{car}_{t},\mathbf{wp}_{i}(x_{i},y_{i})^{closest})$ (6)
 $eth=abs\left(yaw^{car}_{t}-\arctan\left({\frac{y_{i+1}-y_{i}}{x_{i+1}-x_{i}}}\right)\right)$ (7)
-We implement the early-stop condition based on a planned trajectory’s accumulated cross-track error ($xte$) being $<0.1m$ to avoid MPC wasting time on optimizing a trajectory that is already good enough.
-Additionally, to ensure real-time planning during testing, we cap the maximum optimization iterations at 50, which guarantees that the optimization time is less than 0.05 seconds in worst-case scenarios.
-As, 20Hz running rate of MPC is validated by [30] for vehicle path tracking.
-(a) Training tracks
-(b) Zero-shot tracks (complex)
-(c) Zero-shot tracks (simple)
-Figure 3: Tracks tested during training and testing.
-IV-B Baselines and Metrics
-During testing, we evaluate on two metrics: the average MPC optimization time (seconds) per step and the average $xte$ (m) per step.
-We compare our algorithm’s performance against five types of initial guess policies: All Zero: all-zero initial guesses, Previous Solution: initial guesses derived from the MPC solution at the previous step, Ours w/o BC: initial guesses from a warm-start policy trained exclusively with an online training algorithm, without offline training, Ours w/o RL: initial guesses from a warm-start policy trained solely with offline BC, and Ours w/o DAgger: initial guesses from a warm-start policy trained with our two-phase learning algorithm but without the use of DAgger.
+Gradient-based solvers are highly effective when both system dynamics and cost functions are smooth and differentiable.
+However, in our setting, the calculation of xtexte (cross-track error) requires finding the closest point on a reference trajectory to the vehicle’s current position, which introduces discontinuities and non-differentiability into the objective function.
+This motivates the use of gradient-free solvers such as COBYLA [22], which are better suited for non-differentiable optimization problems [27, 28].
+Nonetheless, with high speeds and on tracks with sharp turns, gradient-free MPC solvers struggle to optimize trajectories in real time without a good initial guess, making practical deployment challenging.
+As shown in Section V, the gradient-free MPC solver succeeds only on the simple IMS track but fails on all complex tracks.
+We implement an early-stop condition when the planned trajectory’s accumulated xtexte is below $0.1,m$, preventing MPC wasting time on optimizing a trajectory that is already good enough.
+We also cap MPC iterations at 100 during RL training to avoid solver stagnation and evaluate performance under both 100 and 200 iteration limits during testing for sensitivity analysis.
+IV-A2 RL Details
+The reward at each step of the RL training is shown in Eq.
+(8).
+The first term is the negative MPC optimization time, and the second term is the negative of the xtexte over the planning horizon, $H$.
+This reward design helps optimize $\pi_{\theta}^{warm}$ by minimizing the MPC running time directly while maximizing tracking accuracy.
+$r_{t}=-time_{MPC}-xte(pos_{t}^{car},Traj^{ref})$ (8)
+IV-B Sampling Based MPC
+IV-B1 MPC Details
+Unlike deterministic MPC, MPPI employs a sampling-based stochastic optimization framework to minimize a predefined cost.
+This setup allows us to test whether our warm-start algorithm can also enhance a different control optimizer.
+As MPPI is mainly used for path-planning and obstacle-avoidance tasks [29, 30], we build an obstacle-rich navigation domain for evaluation, as shown in Fig.
+3.
+The training set includes two straight and two curved-road maps, while the testing set contains three unseen vertical and three unseen curved-road maps.
+Each map features randomly generated polygonal obstacles with varying sizes, shapes, and positions, posing diverse planning challenges.
+The vehicle starts at a green dot and must reach the goal region (red dot) without collisions.
+To increase training diversity, each training map offers two initialization points: one on the left and one on the right.
+$J=\sum_{i=t}^{t+H-1}$ $\Big[w_{1}\,throttle_{i}^{2}+w_{2}\,steer_{i}^{2}$ (9)
+$\quad+w_{3}\,(v_{i}-v_{t}^{ref})^{2}$
+$\quad+w_{4}\,C_{obstacle}(x_{i},y_{i})$
+$\quad+w_{5}\,C_{border}(x_{i},y_{i})\Big]$
+$+\quad\sum_{i=t+H-3}^{t+H-1}\Big[w_{6}\,C_{goal}(x_{i},y_{i})\Big]$
+The vehicle dynamics model is identical to that employed in the deterministic MPC in Eq.
+(4), ensuring a fair comparison.
+The cost function for MPPI is shown in Eq.
+(9).
+The first two terms correspond to smoothness penalties to enhance control stability.
+$C_{obstacle}$ penalizes proximity to obstacles, $C_{border}$ penalizes deviations beyond road boundaries to maintain lane centering, and $C_{goal}$ penalizes the final distance to the target region.
+The planning horizon is fixed at 25 steps.
+And number of samples is set to 200 samples to balance efficiency and control quality.
+(a) Training Maps
+(b) Testing Maps
+Figure 3: Training and testing maps in the obstacle-rich navigation domain. Green dots indicate spawn points, red dots represent destinations, and polygons correspond to obstacles.
+IV-B2 RL Details
+The RL reward function is defined in Eq.
+(10).
+It is aligned with the MPPI objective function.
+The only difference is the final term, which encourages the vehicle to reduce its distance to the goal at each timestep.
+Since MPPI operates via parallel trajectory sampling rather than iterative optimization, we exclude the MPC optimization time from the RL reward, as all trajectories are sampled simultaneously on the GPU.
+However, warm-starting remains valuable for onboard systems with limited parallelism.
+A good initial guess allows the controller to achieve comparable performance with fewer samples or higher accuracy with the same number of samples.
+In our evaluation, we fix the number of samples and test whether our method outperforms the baselines in control performance.
+$r_{t}^{mppi}$ $=-w_{1}\,throttle_{t}^{2}-w_{2}\,steer_{t}^{2}$ (10)
+$\quad-w_{3}\,(v_{t}-v_{t}^{ref})^{2}-w_{4}\,C_{obstacle}(x_{t},y_{t})$
+$\quad-w_{5}\,C_{border}(x_{t},y_{t})$
+$\quad-w_{6}\,((C_{goal}(x_{t},y_{t})-(C_{goal}(x_{t-1},y_{t-1}))$
+TABLE I: Deterministic MPC results with a maximum of 100 optimization iterations.
+Values represent averages over three test runs with different random seeds.
+Reported percentages indicate relative improvement compared to the corresponding baseline.
+The asterisk (*) implies statistical significance (paired Wilcoxon signed-rank tests): * $p<0.05$, ** $p<0.01$, *** $p<0.001$, **** $p<0.0001$, n.s.
+means no significance.
+All Zeros &
+Previous Solution &
+Ours w/o BC
+Ours
+w/o RL
+Ours
+w/o DAgger
+Ours
+Optimization
+Time (Second)
+xte
+Optimization
+Time (Second)
+xte
+Optimization
+Time (Second)
+xte
+Optimization
+Time (Second)
+xte
+Training Tracks
+Zandvoort 0.1270 ± 0.0800 0.5929 ± 0.6066 0.1530 ± 0.0542 0.3869 ± 0.2840 0.1087 ± 0.0821 0.3975 ± 0.3084
+Sochi 0.1429 ± 0.0999 0.5657 ± 0.5869 0.1772 ± 0.0620 0.4704 ± 0.5150 0.1128 ± 0.1010 0.3755 ± 0.3075
+Hockenheim Failed 0.1129 ± 0.0912 0.4686 ± 0.4177 0.1450 ± 0.0581 0.4078 ± 0.4665 0.0865 ± 0.0877 0.3548 ± 0.3373
+Training Tracks
+Total
+0.1285 ± 0.0919 0.5450 ± 0.5504 0.1597 ± 0.0600 0.4246 ± 0.4385 0.1033 ± 0.0920 0.3758 ± 0.3179
+Compared with
+Ours w/o RL
+-24.27%
+n.s.
+22.10%
+19.65%
+30.04%
+Zero-shot Tracks
+Nuerburgring 0.1845 ± 0.0669 1.6630 ± 1.5488 0.1679 ± 0.0601 0.3774 ±0.3233 0.1158 ± 0.0967 0.4090 ± 0.3500
+BrandsHatch 0.1163 ± 0.0922 0.5117 ± 0.5484 0.1545 ± 0.0519 0.3463 ± 0.2075 0.0895 ± 0.0837 0.3355 ± 0.2592
+Oschersleben 0.1195 ± 0.0832 0.4181 ± 0.3514 0.1421 ± 0.0582 0.4383 ± 0.4309 0.1116 ± 0.0846 0.3405 ± 0.2395
+Shanghai Failed 0.1845 ± 0.0732 1.3278 ± 1.2934 0.1717 ± 0.0602 0.4242 ± 0.4157 0.1189 ± 0.0977 0.3869 ± 0.2688
+Melbourne 0.1375 ± 0.0963 0.7781 ± 0.9978 0.1767 ± 0.0588 0.4297 ± 0.3378 0.1235 ± 0.0937 0.3921 ± 0.3028
+Catalunya 0.1140 ± 0.0942 0.5561 ± 0.6710 0.1561 ± 0.0571 0.4095 ± 0.3798 0.1019 ± 0.0891 0.3631 ± 0.2863
+Zero-shot Tracks
+Total
+0.1558 ± 0.0871 1.0696 ± 1.2447 0.1629 ± 0.0589 0.4033 ± 0.3539 0.1095 ± 0.0913 0.3707 ± 0.2905
+Compared with
+Ours w/o RL
+-4.57%
+n.s.
+62.30%
+29.75%
+65.34%
+TABLE II: Deterministic MPC results with a maximum of 200 optimization iterations.
+Values represent averages over three test runs with different random seeds.
+Reported percentages indicate relative improvement compared to the corresponding baseline.
+The asterisk (*) implies statistical significance (paired Wilcoxon signed-rank tests): * $p<0.05$, ** $p<0.01$, *** $p<0.001$, **** $p<0.0001$, n.s.
+means no significance.
+All Zeros &
+Previous Solution &
+Ours w/o BC
+Ours
+w/o RL
+Ours
+w/o DAgger
+Ours
+Optimization
+Time (Second)
+xte
+Optimization
+Time (Second)
+xte
+Optimization
+Time (Second)
+xte
+Optimization
+Time (Second)
+xte
+Training Tracks
+Zandvoort 0.2395 ± 0.1700 0.5130 ± 0.5054 0.2498 ± 0.1412 0.3472 ± 0.2664 0.1954 ± 0.1753 0.3501 ± 0.2402
+Sochi 0.2821 ± 0.2079 0.5367 ± 0.5056 0.2836 ± 0.1649 0.3783 ± 0.3697 0.2166 ± 0.2122 0.3621 ± 0.2745
+Hockenheim Failed 0.2072 ± 0.1914 0.4245 ± 0.3663 0.2349 ± 0.1502 0.3754 ± 0.4123 0.1593 ± 0.1815 0.3172 ± 0.2734
+Training Tracks
+Total
+0.2451 ± 0.1935 0.4944 ± 0.4695 0.2578 ± 0.1545 0.3674 ± 0.3546 0.1918 ± 0.1930 0.3441 ± 0.2645
+Compared with
+Ours w/o RL
+-5.19%
+n.s.
+25.70%
+21.77%
+30.41%
+Zero-shot Tracks
+Nuerburgring 0.3086 ± 0.1865 0.5814 ± 0.5536 0.2685 ± 0.1576 0.3359 ±0.3251 0.2358 ± 0.2008 0.4254 ± 0.3867
+BrandsHatch 0.2172 ± 0.1887 0.4122 ± 0.3194 0.2514 ± 0.1421 0.3063 ± 0.1850 0.1584 ± 0.1771 0.3213 ± 0.2282
+Oschersleben 0.2340 ± 0.1754 0.3977 ± 0.3196 0.2467 ± 0.1473 0.3623 ± 0.2838 0.2376 ± 0.1756 0.3893 ± 0.2840
+Shanghai Failed 0.3093 ± 0.1943 0.6666 ± 0.6522 0.2806 ± 0.1551 0.3402 ± 0.3107 0.2150 ± 0.2028 0.3765 ± 0.2549
+Melbourne 0.2747 ± 0.2011 0.6036 ± 0.6359 0.2930 ± 0.1561 0.3868 ± 0.3315 0.2223 ± 0.2016 0.3584 ± 0.2459
+Catalunya 0.2300 ± 0.1947 0.5820 ± 0.6688 0.2663 ± 0.1521 0.3896 ± 0.3693 0.1920 ± 0.1871 0.3704 ± 0.3305
+Zero-shot Tracks
+Total
+0.2663 ± 0.1949 0.5549 ± 0.5710 0.2697 ± 0.1531 0.3543 ± 0.3112 0.2078 ± 0.1921 0.3689 ± 0.2909
+Compared with
+Ours w/o RL
+-1.30%
+n.s.
+36.16%
+21.97%
+33.52%
+IV-C Baselines and Metrics
+For deterministic MPC, we evaluate performance using two metrics: (1) the average optimization time per step (seconds) and (2) the average xtexte (meters) per step.
+For sampling-based MPC, we employ three metrics: (1) steering standard deviation, to assess control stability; (2) average number of steps required to reach the goal, as a proxy for path efficiency; and (3) number of obstacle collisions, to evaluate safety and avoidance capability.
+During testing, we compare the performance of our algorithm against five types of initial-guess policies:
+  • All Zero: all-zero initial guesses.
+  • Previous Solution: initial guesses derived from the MPC solution at the previous step.
+  • Ours w/o BC: initial guesses from a warm-start policy trained exclusively with an online training algorithm, without offline training.
+  • Ours w/o RL: initial guesses from a warm-start policy trained solely with offline BC.
+  • Ours w/o DAgger: initial guesses from a warm-start policy trained with our two-phase learning algorithm but without the use of DAgger.
 To ensure a fair comparison between the Ours w/o BC and warm-start policy trained by our two-phase learning algorithm, we extend the training time of Ours w/o BC by an additional hour to account for the data collection time in the offline BC phase.
-IV-C Experiment Results
-We perform testing on both the training tracks (Figure 3(a)) and the challenging zero-shot tracks (Figure 3(b)).
-The results are shown in Figure 4 and Figure 5, respectively.
+TABLE III: Sampling Based MPC Results. Values represent averages over three test runs with different random seeds. Numbers in parentheses denote the standard error.
+All Zeros Previous Solution
+Ours
+w/o BC
+Ours
+w/o RL
+Ours
+w/o DAgger
+Ours
+Map
+Name
+Steer
+Std
+Number
+of Steps
+Number
+of Hits
+Steer
+Std
+Number
+of Steps
+Number
+of Hits
+Steer
+Std
+Number
+of Steps
+Number
+of Hits
+Steer
+Std
+Number
+of Steps
+Number
+of Hits
+Steer
+Std
+Number
+of Steps
+Number
+of Hits
+Steer
+Std
+Number
+of Steps
+Number
+of Hits
+Training Maps
+Vertical_1 0.6647
+287.3333
+± 22.8356
+2.0000
+± 1.4142
+0.6306
+251.8333
+± 12.0568
+0.0000
+± 0.0000
+0.4881
+258.0000
+± 19.7281
+1.0000
+± 1.0954
+0.4747
+278.1667
+± 16.7978
+0.6667
+± 0.8165
+0.4712
+245.1667
+± 20.8750
+0.8333
+± 1.6021
+0.4689
+249.3333
+± 22.2860
+0.0000
+± 0.0000
+Vertical_2 0.6491
+289.0000
+± 45.4093
+0.6667
+± 1.2111
+0.6393
+273.6667
+± 11.5873
+0.0000
+± 0.0000
+0.4923
+264.0000
+± 23.0304
+0.8333
+± 0.9832
+0.4637
+279.8333
+± 24.7258
+0.0000
+± 0.0000
+0.4810
+265.6667
+± 19.2942
+0.0000
+± 0.0000
+0.4716
+256.5000
+± 26.8384
+0.1667
+± 0.4082
+Curved_1 0.6290
+369.8333
+± 37.0751
+1.5000
+± 3.6742
+0.6009
+360.3333
+± 13.4263
+0.0000
+± 0.0000
+0.4773
+376.3333
+± 84.5167
+0.0000
+± 0.0000
+0.4748
+354.6667
+± 36.9035
+0.0000
+± 0.0000
+0.4322
+334.6667
+± 4.1793
+0.0000
+± 0.0000
+0.4256
+335.0000
+± 5.6569
+0.0000
+± 0.0000
+Curved_2 0.6366
+422.6667
+± 65.9019
+9.3333
+± 17.5803
+0.6369
+401.3333
+± 26.1738
+0.0000
+± 0.0000
+0.5064
+427.6667
+± 60.8857
+1.0000
+± 2.4495
+0.4879
+460.5000
+± 81.4389
+0.0000
+± 0.0000
+0.4812
+375.6667
+± 24.7117
+1.3333
+± 2.0656
+0.4800
+376.1667
+± 17.6682
+0.0000
+± 0.0000
+Training Map
+Total
+0.6433
+342.2083
+± 72.2255
+3.3750
+± 9.1358
+0.6266
+321.7917
+± 64.5587
+0.0000
+± 0.0000
+0.4922
+331.5000
+± 89.9633
+0.7083
+± 1.3981
+0.4775
+343.2917
+± 87.7870
+0.1667
+± 0.4815
+0.4663
+305.2917
+± 56.4088
+0.5417
+± 1.3507
+0.4619
+304.2500
+± 57.5917
+0.0417
+± 0.2041
+Testing Maps
+Vertical_1 0.6425
+285.6667
+± 107.7698
+9.6667
+± 7.7675
+0.6228
+288.3333
+± 97.6183
+0.0000
+± 0.0000
+0.4667
+225.0000
+± 1.0000
+0.0000
+± 0.0000
+0.4982
+355.6667
+± 198.4574
+0.0000
+± 0.0000
+0.4697
+230.0000
+± 21.7945
+0.3333
+± 0.5774
+0.4579
+220.6667
+± 4.5093
+0.0000
+± 0.0000
+Vertical_2 0.6416
+266.3333
+± 35.5575
+2.0000
+± 2.6458
+0.6373
+231.6667
+± 3.0551
+0.0000
+± 0.0000
+0.5068
+235.6667
+± 14.2244
+1.0000
+± 1.0000
+0.4709
+225.3333
+± 4.1633
+0.6667
+± 1.1547
+0.4592
+217.6667
+± 3.2146
+0.0000
+± 0.0000
+0.4421
+219.3333
+± 14.4338
+0.0000
+± 0.0000
+Vertical_3 0.6298
+312.0000
+± 106.2262
+3.3333
+± 4.9329
+0.6280
+263.6667
+± 38.5011
+0.0000
+± 0.0000
+0.4332
+220.3333
+± 1.1547
+0.0000
+± 0.0000
+0.4646
+258.0000
+± 4.3589
+0.0000
+± 0.0000
+0.4685
+257.3333
+± 1.5275
+0.0000
+± 0.0000
+0.4621
+239.0000
+± 25.1197
+0.0000
+± 0.0000
+Curved_1 0.6251
+348.0000
+± 2.6458
+9.6667
+± 12.6623
+0.5929
+328.6667
+± 3.0551
+0.0000
+± 0.0000
+0.4457
+323.6667
+± 3.5119
+0.0000
+± 0.0000
+0.4385
+328.0000
+± 10.5357
+0.0000
+± 0.0000
+0.4267
+317.6667
+± 0.5774
+0.0000
+± 0.0000
+0.4227
+315.6667
+± 1.1547
+0.0000
+± 0.0000
+Curved_2 0.5890
+331.0000
+± 9.1652
+0.3333
+± 0.5774
+0.5720
+319.6667
+± 0.5774
+0.0000
+± 0.0000
+0.4163
+314.0000
+± 1.7321
+0.0000
+± 0.0000
+0.4060
+310.6667
+± 1.1547
+0.0000
+± 0.0000
+0.4130
+308.6667
+± 2.0817
+0.0000
+± 0.0000
+0.4120
+310.0000
+± 4.5826
+0.0000
+± 0.0000
+Curved_3 0.6416
+418.0000
+± 83.5165
+3.3333
+± 3.5119
+0.5841
+339.3333
+± 8.1445
+0.0000
+± 0.0000
+0.4654
+340.6667
+± 26.3881
+0.0000
+± 0.0000
+0.4975
+387.6667
+± 69.8164
+1.3333
+± 2.3094
+0.4446
+324.3333
+± 5.6862
+0.0000
+± 0.0000
+0.4330
+326.0000
+± 7.2111
+0.0000
+± 0.0000
+Testing Map
+Total
+0.6284
+326.8333
+± 78.7821
+4.7222
+± 6.7196
+0.6050
+295.2222
+± 53.3728
+0.0000
+± 0.0000
+0.4550
+276.5556
+± 52.8597
+0.1667
+± 0.5145
+0.4713
+310.8889
+± 91.9545
+0.3333
+± 1.0290
+0.4467
+275.9444
+± 44.7667
+0.0556
+± 0.2357
+0.4372
+271.7778
+± 48.6047
+0.0000
+± 0.0000
+TABLE IV: Relative improvement between different methods in sampling-based MPC and statistical test results.
+The reported percentages indicate relative improvement.
+The asterisk (*) implies statistical significance (paired Wilcoxon signed-rank tests): * $p<0.05$, ** $p<0.01$, *** $p<0.001$, **** $p<0.0001$, n.s.
+means no significance.
+Previous Solution
+Ours
+w/o BC
+Ours
+w/o RL
+Ours
+w/o DAgger
+Ours
+Steer
+Std
+Number
+of Steps
+Number
+of Hits
+Steer
+Std
+Number
+of Steps
+Number
+of Hits
+Steer
+Std
+Number
+of Steps
+Number
+of Hits
+Steer
+Std
+Number
+of Steps
+Number
+of Hits
+Steer
+Std
+Number
+of Steps
+Number
+of Hits
+Training Maps
+Compared with
+All Zeros
+2.60%
+5.97%
+100%
+23.49%
+3.13%
+n.s.
+79.01%
+25.77%
+-0.32%
+n.s.
+95.06%
+27.51%
+10.79%
+83.95%
+28.20%
+11.09%
+98.76%
+Compared with
+Previous Solution
+21.45%
+-3.02%
+n.s.
+23.80%
+-6.68%
+n.s.
+25.58%
+5.13%
+26.27%
+5.46%
+Compared with
+Ours w/o BC
+2.99%
+-3.56%
+n.s.
+76.46%
+5.26%
+7.91%
+23.52%
+n.s.
+6.16%
+8.22%
+94.11%
+Compared with
+Ours w/o RL
+2.35%
+n.s.
+11.07%
+-224.96%
+n.s.
+3.27%
+n.s.
+11.37%
+74.99%
+n.s.
+Compared with
+Ours w/o DAgger
+0.95%
+n.s.
+0.34%
+n.s.
+92.30%
+Testing Maps
+Compared with
+All Zeros
+3.72%
+9.67%
+100.00%
+27.59%
+15.38%
+96.47%
+25.00%
+4.88%
+92.94%
+28.91%
+15.57%
+98.82%
+30.43%
+16.85%
+100.00%
+Compared with
+Previous Solution
+24.79%
+6.32%
+22.10%
+-5.31%
+n.s.
+26.17%
+6.53%
+27.74%
+7.94%
+0.00%
+Compared with
+Ours w/o BC
+-3.58%
+n.s.
+-12.41%
+n.s.
+-99.94%
+n.s.
+1.82%
+n.s.
+0.22%
+n.s.
+66.65%
+n.s.
+3.91%
+1.73%
+100.00%
+n.s.
+Compared with
+Ours w/o RL
+5.21%
+n.s.
+11.24%
+83.32%
+n.s.
+7.22%
+12.58%
+100.00%
+n.s.
+Compared with
+Ours w/o DAgger
+2.13%
+1.51%
+n.s.
+100.00%
+n.s.
+V Experiment Results
+In this section we design experiments to test the following research questions in both deterministic MPC and sampling-based MPC problem:
+  • RQ1 Does our warm-start policy reduce solver runtime in deterministic MPC problem?
+  • RQ2 Does our warm-start policy improve tracking performance in deterministic MPC problem?
+  • RQ3 Does our warm-start policy generalize to unseen tracks in deterministic MPC problem?
+  • RQ4 Does our warm-start policy generalize from deterministic MPC to sampling-based MPC problem?
+V-A Deterministic MPC
+We perform testing on both the training tracks (Fig.
+2(a)) and the challenging zero-shot tracks (Fig.
+2(b)).
+The results are shown in Table I.
 On both the training and zero-shot tracks, employing a warm-start policy trained through either offline learning or a combination of offline and online fine-tuning significantly reduces MPC optimization time and improves tracking accuracy.
-Besides, the warm-start policy trained with both offline and online fine-tuning achieves better MPC optimization time and tracking accuracy compared to the policy trained solely through offline BC, demonstrating the capability of our online fine-tuning algorithm in addressing the suboptimality covariance shift problem caused by BC.
-[…]
-The results also reveal that initialize an MPC with either All Zeros or Previous Solution fails to complete laps on both the training and zero-shot tracks.
+Besides, the warm-start policy trained with both offline and online fine-tuning achieves better MPC optimization time and tracking accuracy compared to the policy trained solely via offline BC, demonstrating the capability of our online fine-tuning algorithm in addressing the suboptimality and covariance shift problem in BC.
+On the training tracks, compared to Ours w/o RL, the proposed method improves optimization time by $19.95\%$ and tracking accuracy by $30.60\%$.
+When compared to Ours w/o RL, Ours w/o DAgger achieves a $19.12\%$ improvement in tracking accuracy; however, the optimization time degrades by $22.08\%$, highlighting the necessity of incorporating the DAgger training loss during RL training.
+On the zero-shot tracks, compared to Ours w/o RL, the proposed method improves optimization time by $21.63\%$ and tracking accuracy by $34.12\%$.
+Similarly, Ours w/o DAgger achieves a $30.35\%$ improvement in tracking accuracy, but optimization time degrades by $18.00\%$.
+This degradation pattern aligns with that on the training tracks, further highlighting the importance of integrating the DAgger training loss during RL training.
+Figure 4: Visualization of optimized trajectories for our method and the baselines on a segment of the Nürburgring map.
+The results also reveal that initializing an MPC with either All Zeros or Previous Solution fails to complete laps on both the training and zero-shot tracks.
 The vehicle consistently deviates from the lane when facing any sharp turns.
 This limitation arises because the MPC often requires more iterations to optimize the control solution as the vehicle approaches the curves of the track.
-Since the real-time MPC only optimizes the solution for a maximum of 50 iterations at each step, the returned solution lacks the optimization necessary to guide the vehicle through sharp turns effectively.
+Since the real-time MPC only optimizes the solution for a maximum of 100 iterations at each step, the returned solution lacks the optimization necessary to guide the vehicle through sharp turns effectively.
 This underscores the necessity of a well-informed initial guess to minimize the number of optimization iterations required.
-Figure 4: Experiment results on training tracks. The asterisk (*) implies statistical significance: * $p<0.05$, ** $p<0.01$, *** $p<0.001$, n.s. means no significant.
-Figure 5: Experiment results on the six zero-shot testing tracks. The asterisk (*) implies statistical significance: * $p<0.05$, ** $p<0.01$, *** $p<0.001$, n.s. means no significant.
-Furthermore, the Ours w/o RL algorithm also fails on both training and zero-shot domains.
-This is because the absence of offline BC requires the algorithm to explore the solution space from scratch, causing the online learning algorithm to take significantly longer to converge to a satisfactory solution.
-Given the same amount of training time, the quality of the policy learned by Ours w/o RL lags behind that of the policy learned through the two-phase training algorithm.
-This demonstrates the effectiveness of offline BC in accelerating online training.
-To further inspect the relative performance of our approach when our baselines performed better, i.e.
-on simpler tracks, we conducted further testing on six warm-start policies on a simpler zero-shot track, IMS, as depicted in Figure 3(c),   similar to the path-tracking scenario in [31].
-The results are summarized in Figure 6.
-Our method demonstrates a significant reduction in MPC optimization time while maintaining high tracking accuracy, showcasing its effectiveness in enhancing MPC performance.
-Figure 6: Experiment results on the simple IMS track.
+We also relaxed the maximum MPC iterations to 200 and repeated the test.
+The results in Table II show that our warm-start policy again achieves the greatest improvement in optimization time and the largest performance gains over the baselines, further supporting RQ1, RQ2, and RQ3.
+A visualization of the optimized trajectories for our method and the baselines on a segment of the Nürburgring map is presented in Fig.
+4 (MPC optimization iterations=100).
+The figure illustrates that using our method as a warm-start strategy enables the MPC to track the reference path more accurately by tracking the sharp more accurately.
+V-B Sampling-Based MPC
+We evaluate performance on the training maps (Fig.
+3(a)) and the zero-shot maps (Fig.
+3(b)), with results summarized in Table III.
+Consistent with the findings for deterministic MPC, our proposed framework delivers the most significant performance improvements for the MPPI controller compared to all baselines.
+Notably, the improvements are more pronounced in the zero-shot environments, demonstrating the strong generalizability of the proposed method.
+V-C Findings
 In summary, our empirical results support that:
-  • MPC with warm-start policy trained by either offline learning or a combination of offline and online fine-tuning significantly outperforms those with all-zero or previous solution-derived guesses, crucial for navigating sharp turns.
-  • The offline training algorithm significantly warm-starts the online training algorithm and make it take much less training time to get a satisfying performance.
-  • The warm-start policy trained with our two-phase learning framework reduces MPC optimization time and $xte$ on both training and zero-shot tracks, showing the capability of our online fine-tuning algorithm in addressing the suboptimality problem.
-  • The two-phase training algorithm outperforms in zero-shot scenarios, demonstrating superior generalizability.
-V LIMITATIONS AND FUTURE WORKS
-While our proposed two-phase learning framework shows promising results in expediting optimization processes and enhancing control performance for robot control tasks, there are still limitations and avenues for future research that merit consideration.
-Future work could expand into more complex domains, such as robot arm manipulation or quadcopter control, to explore the algorithm’s efficacy in higher-dimensional spaces.
-Additionally, future research could also involve utilizing our proposed algorithm to expedite other iterative control optimization techniques like MPPI.
-Although our algorithm primarily interfaces with MPC, it is versatile enough to serve as a warm start for other iterative control optimization methods like Model Predictive Path Integral (MPPI) control [32].
-MPPI typically employs stochastic optimization, initializing with a guess of the control sequence distribution, sampling sequences, and iteratively refining the distribution to minimize control cost.
-Our framework can be trained to predict the trajectory Gaussian distribution mean, enhancing MPPI’s performance by offering a superior initial distribution.
-This reduces the iteration count needed for optimization, aligning with early stopping criteria.
-While MPC suits systems with complex dynamics and constraints, we choose it as our focus due to its representativeness in trajectory optimization methods.
-VI CONCLUSIONS
+  • Our proposed framework, trained offline or with online fine-tuning, outperform other non-learning-based baselines.
+  • The two-phase training algorithm generalizes better on zero-shot tracks, achieving superior zero-shot performance.
+  • Offline training effectively accelerates online fine-tuning, reducing training time to achieve strong performance.
+VI LIMITATIONS AND FUTURE WORKS
+While our proposed two-phase learning framework shows promising results in expediting optimization processes and enhancing control performance for robot control tasks, it also has several limitations for future research that merit consideration.
+First, our experiments assume access to perfect state information.
+Second, the current evaluation is limited to vehicle control.
+Future work could expand into higher-DOF robotic platforms such as manipulators or aerial robots.
+VII CONCLUSIONS
 In this paper, we introduce a novel approach to accelerate MPC optimization by learning a warm-start policy.
 Our two-phase framework combines offline BC and online fine-tuning to provide better initial guesses for the MPC solver.
-Experiments on training and zero-shot tracks demonstrate the effectiveness of our approach in reducing optimization time and improving path tracking accuracy.
+Experiments on training and zero-shot tracks demonstrate the effectiveness of our approach in reducing optimization time without degrading MPC’s performance.
 This integration of learning with MPC enhances the efficiency and applicability of trajectory optimization in dynamic systems.